### v2

# Human-in-the-Loop Causal Discovery under Latent Confounding using Ancestral GFlowNets

Abstract
Structure learning is the crux of causal inference. Notably, causal discovery (CD) algorithms are brittle when data is scarce, possibly inferring imprecise causal relations that contradict expert knowledge — especially when considering latent confounders. To aggravate the issue, most CD methods do not provide uncertainty estimates, making it hard for users to interpret results and improve the inference process. Surprisingly, while CD is a human-centered affair, no works have focused on building methods that both 1) output uncertainty estimates that can be verified by experts and 2) interact with those experts to iteratively refine CD. To solve these issues, we start by proposing to sample (causal) ancestral graphs proportionally to a belief distribution based on a score function, such as the Bayesian information criterion (BIC), using generative flow networks. Then, we leverage the diversity in candidate graphs and introduce an optimal experimental design to iteratively probe the expert about the relations among variables, effectively reducing the uncertainty of our belief over ancestral graphs. Finally, we update our samples to incorporate human feedback via importance sampling. Importantly, our method does not require causal sufficiency (i.e., unobserved confounders may exist). Experiments with synthetic observational data show that our method can accurately sample from distributions over ancestral graphs and that we can greatly improve inference quality with human aid.
1 Introduction
Drawing conclusions about cause-and-effect relationships presents a fundamental challenge in various scientific fields and significantly impacts decision-making across diverse domains Pearl (2000). The importance of having structural knowledge, often encoded as a causal diagram, for conducting causal inferences is widely recognized, a concept made prominent by Cartwright (1989)’s dictum: “no causes in, no causes out”. When there is no objective knowledge to fully specify a causal diagram, causal discovery (CD) tools are instrumental in partially uncovering causal relationships among variables from, for example, observational data. Formally, let $\mathbf{V}=\{V_{1},V_{2},\dots,V_{n}\}$ be a set of $n$ observed variables and $\mathcal{D}$ be a dataset containing $|\mathcal{D}|=m$ samples for each $V_{i}\in\mathbf{V}$. A CD algorithm takes $\mathcal{D}$ as input and typically returns a single graph $\mathcal{G}=(\mathbf{V},\mathbf{E})$ with well-defined causal semantics, in which each node in $\mathbf{V}$ represents a variable $V_{i}\in\mathbf{V}$ and each edge in $\mathbf{E}$ encodes the possible underlying (causal/confounding) mechanisms compatible with $\mathcal{D}$.
Figure 1: Human-in-the-loop probabilistic CD. We first train an AGFN to fit a data-informed belief over AGs. Then, we iteratively refine it by 1) questioning (Q) experts on the relation between a highly informative pair of variables and 2) updating the belief given the potentially noisy answers (A). The histograms on top of the edges show marginals over edge types (green denotes ground truth). Notably, our belief increasingly concentrates on the true AG, $1\rightarrow 2\leftrightarrow 3$.
This work focuses on recovering the structure of the underlying causal diagram when unobserved confounders may be at play. We propose to address this task by not only leveraging observational data but also by accounting for potentially noisy pieces of expert knowledge, otherwise unavailable as data. Throughout this work, we consider ancestral graphs (AGs) as surrogates for causal diagrams. AGs are particularly convenient since they encode latent confounding without explicitly invoking unobserved variables. Moreover, AGs capture all conditional independencies and ancestral relations among observed variables $\mathbf{V}$, as entailed by a causal diagram (Richardson and Spirtes, 2002).
In the realm of CD from solely observational data, algorithms aim to construct a compact representation of the joint observational distribution $P(\mathbf{V})$, which implies a factorization as a product of conditional probabilities. Notably, multiple models may entail the same conditional independencies; in such cases, they are denoted as Markov-equivalent. As a result, these algorithms can only reconstruct the class of Markov-equivalent models (AGs), denoted as the Markov Equivalence Class (MEC) and typically represented by a Partial Ancestral Graph (PAG). Importantly, CD beyond the MEC by leveraging domain knowledge presents a critical challenge. Notably, there is no proper characterization of an equivalence class that accounts for knowledge stemming from both humans and data (Wang et al., 2022).
There is a variety of algorithms for CD from observational data, primarily categorized into constraint- and score-based methods. The former uses (in)dependence constraints derived via conditional independence tests to directly construct a PAG representing the MEC. The latter uses a goodness-of-fit score to navigate the space of AGs, selecting an optimum as a representative for the MEC.
Nonetheless, methods within both paradigms suffer from unreliability when data is scarce. Specifically, for the majority of the CD algorithms, formal assurances that the inferred MEC accurately represents the true causal model heavily rely on the so-called faithfulness assumption, which posits that all conditional independencies satisfied by $P(\mathbf{V})$ are entailed by the true causal model (Zhang and Spirtes, 2016). However, this presents a critical challenge in real-world scenarios, as violations of the faithfulness assumption become more prominent when relying on $P(\mathbf{V})$ estimated from limited data Uhler et al. (2012); Andersen (2013); Marx et al. (2021). For constraint-based methods, hypothesis tests may lack the statistical power to detect conditional independencies accurately. These errors may propagate and trigger a chain reaction of erroneous orientations Zhang and Spirtes (2008); Zhalama et al. (2017b); Ng et al. (2021). For score-based methods, although score functions directly measure goodness-of-fit on observational data, small sample sizes can significantly skew the estimates for the population parameters. Consequently, structures deemed score-optimal may not necessarily align with the ground-truth MEC  Ogarrio et al. (2016). A major concern is that the overwhelming majority of CD algorithms produce a single representation of the MEC as output, without quantifying the uncertainty that arises during the learning process Claassen and Heskes (2012); Jabbari et al. (2017). This poses a significant challenge for experts, as it hinders their ability to validate the algorithm’s outcome or gain insights into potential venues for improving inference quality.
To alleviate the lack of uncertainty quantification in CD, we propose sampling AGs from a distribution defined using a score function, which places best-scoring AGs around the mode by design. This effectively provides end-users with samples that reflect the epistemic uncertainty inherent in CD, thus allowing their propagation through downstream causal inferences. In particular, we sample from our belief using Generative Flow Networks (GFlowNets; Bengio et al., 2021a, b), which are generative models known for sampling diverse modes while avoiding the mixing time problem of MCMC methods, and not requiring handcrafted proposals nor accept-reject steps (Bengio et al., 2021a).
Acknowledging the low-data regime as CD’s primary challenge, we also propose actively integrating human feedback in the inferential process. This involves modeling user knowledge on the existence and nature (confounding/ancestral) of the relations and using it to weigh our beliefs over AGs. During our interactions with experts, we probe them about the relation of the pair of variables that optimizes a utility/acquisition function, for which we propose the negative expected cross-entropy between our prior and updated beliefs. Unlike prior strategies, our acquisition avoids the need to estimate the normalizing constant and predictive distribution of our updated belief, as needed for information gain and mutual information, respectively. Notably, we use importance sampling (Marshall, 1954; Geweke, 1989) to update our initial belief with the human feedback, which avoids retraining GFlowNets after each human interaction.
While incorporating expert knowledge into CD has been a long-standing goal (Meek, 1995a; Chen et al., 2016; Li and Beek, 2018; Wang et al., 2022), existing works either rely on strong assumptions (e.g., causal sufficiency) or assume the knowledge is noiseless, aligned with the ground truth  (Andrews, 2020; Wang et al., 2022). Importantly, our work introduces the first iterative CD framework for AGs involving a human in the loop and accommodating potentially noisy feedback, as depicted in Figure 1.
To validate our approach, we conduct experiments using the BIC score, for linear Gaussian causal models. Specifically, we assess: i) our ability to sample from score-based beliefs over AGs, ii) how our samples compare to samples from bootstrapped versions of state-of-the-art (SOTA) methods, and iii) the efficacy of our active knowledge elicitation framework using simulated humans. We observe that our method, Ancestral GFlowNet (AGFN), i) accurately samples from our beliefs over AGs; ii) consistently includes AGs with low structural error among its top-scored samples; and iii) is able to greatly improve performance metrics (i.e., SHD and BIC) when incorporating human in the loop.
In summary, the contributions of our work are:
  1. 1.
    We leverage GFlowNets to introduce AGFN, the first CD algorithm for scenarios under latent confounding that employs fully-probabilistic inference on AGs;
  2. 2.
    We show AGFN accurately learns distributions over AGs, effectively capturing epistemic uncertainty.
  3. 3.
    We propose an experimental design to query potentially noisy expert insights on relationships among pairs of variables that lead to optimal uncertainty reduction.
  4. 4.
    We show how to incorporate expert feedback into AGFN without retraining them from scratch.
2 Background
This section introduces the relevant notation and concepts. We use uppercase letters $V$ to represent a random variable or node in a graph, and boldface uppercase letters $\mathbf{V}$ to represent matrices or sets of random variables or nodes.
Ancestral graphs. Under the assumption of no selection bias, an ancestral graph (AG) $\mathcal{G}$ over $\mathbf{V}$ is a directed graph comprising directed ($\rightarrow$) and bidirected ($\leftrightarrow$) edges Richardson and Spirtes (2002); Zhang (2007). In any directed graph, if a sequence of directed edges, $V_{i}\rightarrow\cdots\rightarrow V_{j}$, connects two nodes $V_{i}$ and $V_{j}$, we refer to this sequence as a directed path. In this case, we also say that $V_{i}$ is an ancestor of $V_{j}$ and denote this relation as $V_{i}\in An(V_{j})$. By definition, any AG $\mathcal{G}$ must further satisfy the following:
  1. 1.
    there is no directed cycle, i.e., if $V_{i}\rightarrow V_{j}$ is in $\mathcal{G}$, then $V_{j}\not\in An(V_{i})$; and
  2. 2.
    there is no almost directed cycle, i.e., if $V_{i}\leftrightarrow V_{j}$ is in $\mathcal{G}$, then $V_{j}\not\in An(V_{i})$ and $V_{i}\not\in An(V_{j})$.
As a probabilistic model, the nodes in an AG represent random variables, directed edges represent ancestral (causal) relationships, and bidirected edges represent associations solely due to latent confounding. For a complete characterization of AGs, refer to Richardson and Spirtes (2002).
Data generating model. We assume that the data-generating model corresponds to a linear Gaussian structural causal model (SCM) (Pearl, 2000) defined by a 4-tuple $\mathcal{M}=\langle\mathbf{V},\mathbf{U},\mathcal{F},P(\mathbf{U})\rangle$, in which $\mathbf{V}=\{V_{1},V_{2},\dots,V_{n}\}$ is a set of $n$ observed random variables and $\mathbf{U}=\{U_{1},U_{2},\dots,U_{n}\}$ is the set of unobserved random variables. Further, let $Pa_{i}\subseteq\mathbf{V}\setminus\{V_{i}\}$ be the set of observed causes (parents) of $V_{i}$, and $U_{i}$ be the set of unobserved causes of $V_{i}$. Then, each structural equation $f_{i}\in\mathcal{F}$ is defined as:
$V_{i}=\sum_{j:V_{j}\in Pa_{i}}\beta_{ij}V_{j}+U_{i}$ (1)
with $P(\mathbf{U})$ being a multivariate Gaussian distribution with zero mean and a not necessarily identity covariance matrix $\boldsymbol{\Omega}=(\omega_{ij})_{1\leq i,j\leq n}$ — the error terms $\{U_{i}\}$ are not necessarily mutually independent, implying that the system can be semi-Markovian (i.e., latent confounding may be present).
Consider a lower triangular matrix of structure coefficients $\mathbf{B}=(\beta_{ij})_{1\leq i,j\leq n}$ such that $(\mathbf{I}-\mathbf{B})$ is invertible, and $\beta_{ij}\neq 0$ only if $V_{j}\in Pa_{i}$. Then, the set of structural equations is given in matrix form by
$\mathbf{V}=\mathbf{B}\mathbf{V}+\mathbf{U}\implies\mathbf{V}=(\mathbf{I}-\mathbf{B})^{-1}\mathbf{U}.$ (2)
The class of all linear Gaussian SCMs parametrized as
$\mathcal{N}_{\mathcal{M}}=\{\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma})|\boldsymbol{\Sigma}=(\mathbf{I}-\mathbf{B})^{-1}\boldsymbol{\Omega}(\mathbf{I}-\mathbf{B})^{-\top}\}$ (3)
is represented by an AG in which, for every $i\neq j$, there is a directed edge $V_{j}\rightarrow V_{i}$ if $\beta_{ij}\neq 0$ and a bidirected edge $V_{j}\leftrightarrow V_{i}$ if $\omega_{ij}\neq 0$ (Richardson and Spirtes, 2002).
[…]
Figure 2: Illustration of the generative process of AGs $\{\mathcal{G}_{5},\mathcal{G}_{6},\mathcal{G}_{7}\}$ using GFlowNets. Starting with an empty graph $\mathcal{G}_{0}$, we add edges between variables $\{X_{1},X_{2},X_{3}\}$ according to the action-policy $\pi_{F}$. Solid edges trace trajectories leading to sampled graphs. Dashed lines represent non-realized transitions to terminal state $\square$.
GFlowNets. Generative Flow Networks (GFlowNet; Bengio et al., 2021a, b) are generative models designed to sample from a finite domain $\mathcal{X}$ proportionally to some reward function $R:\mathcal{X}\rightarrow\mathbb{R}_{+}$, which may be parametrized using neural networks. In this work, we define $R$ as a strictly decreasing transformation of the BIC (more details in section 3). GFlowNets also assume there is a compositional nature to the elements $x\in\mathcal{X}$, meaning that they can be built by iteratively acting to modify a base object (i.e., an initial state). For instance, graphs can be built by adding edges to a node skeleton (Deleu et al., 2022) or molecules by adding atoms to an initial structure (Bengio et al., 2021a).
The generative process follows a trajectory of states $s\in\mathcal{S}$ guided by a transition probability $\pi_{F}:\mathcal{S}^{2}\rightarrow[0,1]$. In turn, $\pi_{F}$ is proportional to a foward flow function $F_{\theta}:\mathcal{S}^{2}\rightarrow\mathbb{R}_{+}$, which is parameterized by a neural network $\theta$. Let $\text{Pa}(s^{\prime})$ ($\text{Ch}(s^{\prime})$) be the set of all states which can transition into (directly reached from) $s^{\prime}$. Then, $\pi_{F}$ is defined as
$\pi_{F}(s^{\prime}|s)=\frac{F_{\theta}(s\rightarrow s^{\prime})}{\sum_{{s^{\prime}\in\text{Ch}(s)}}F_{\theta}(s\rightarrow s^{\prime})}.$ (4)
The support $\mathcal{X}$ of $R$ is contained within $\mathcal{S}$. There are also two special states in $\mathcal{S}$: an initial state $s_{0}$ and a terminal state $s_{f}$. We start with the initial state $s_{0}$ and transform it to a new valid state $s$ with probability $\pi_{F}(s|s_{0};\theta)$. We keep iterating this procedure until reaching $s_{f}$. States $s$ valid as final samples ($s\in\mathcal{X}$) are known as terminating states and have a positive probability for the transition $s\rightarrow s_{f}$. Figure 2 illustrates this process with $\mathcal{X}$ being the space of AGs. Crucially, the same parameterization $\theta$ is used for all transition probabilities $\pi_{F}(\cdot|s;\theta)$ given any departing state $s$, allowing for generalization to states never visited during training.
As the GFlowNet framework requires that no sequence of actions leads to a loop, we represent the space of possible action sequences by a pointed Directed Acyclic Graph (DAG) (Bengio et al., 2021b). The generation of any sample $x\in\mathcal{X}$ follows a trajectory $\tau=(s_{0},s_{1},\ldots,s_{T}=x,s_{f})\in\mathcal{S}^{T+2}$ for a $T\geq 0$. Different trajectories may lead to the same sample $x$. To ensure we sample proportionally to $R$, we search for a GFlowNet that satisfies the flow-matching condition, i.e., $\forall s^{\prime}\in\mathcal{S}$:
$\sum_{s\in\text{Pa}(s^{\prime})}F_{\theta}(s\rightarrow s^{\prime})=R(s^{\prime})+\sum_{s^{\prime\prime}\in\text{Ch}(s^{\prime})}F_{\theta}(s^{\prime}\rightarrow s^{\prime\prime}).$ (5)
Equation 5 implies the flow that enters $s^{\prime}$ equals the flow leaving $s^{\prime}$, except for some flow $R(s^{\prime})$ leaking from $s^{\prime}$ into $s_{f}$. We let $R(s)=0$ for $s\notin\mathcal{X}$. Eventually, it may be that all states $s$ are valid candidates, i.e., $\mathcal{S}=\mathcal{X}\cup\{s_{f}\}$. If so, each of eq. 5’s solutions satisfies a detailed-balance condition,
$\frac{R(s)F_{\theta}(s\rightarrow s^{\prime})F_{\theta}(s^{\prime}\rightarrow s_{f})}{F_{\theta}(s\rightarrow s_{f})}=R(s^{\prime})F_{B,\theta}(s^{\prime}\rightarrow s),$ (6)
for a parametrized backward flow $F_{B,\theta}\colon\mathcal{S}^{2}\rightarrow\mathbb{R}_{+}$ Deleu et al. (2022). In practice, we enforce eq. 6 by minimizing
$\mathcal{L}(\theta)\!=\!\underset{s\rightarrow s^{\prime}}{\mathbb{E}}\!\left[\!\left(\!\log\frac{R(s^{\prime})\pi_{F_{B}}(s|s^{\prime};\theta)\pi_{F}(s_{f}|s;\theta)}{R(s)\pi_{F}(s^{\prime}|s;\theta)\pi_{F}(s_{f}|s^{\prime};\theta)}\!\right)^{2}\!\right].$ (7)
3 Ancestral GFlowNets
We propose AGFN, a GFlowNet-based method for sampling AGs using a score function. Specifically, AGFN encompasses a GFlowNet with the following characteristics:
  1. 1.
    Each trajectory state is a valid AG $\mathcal{G}_{t}$.
  2. 2.
    A terminating state’s reward $R(\mathcal{G_{T}})$ is a score-based potential suitable for CD of AGs.
  3. 3.
    A well-trained AGFN samples AGs with frequencies proportional to their rewards and with the best-scoring AG being, by design, the mode.
[…]
Action constraints. To ensure AGFN only samples AGs, we mask out actions that would lead to paths forming cycles or almost cycles. To achieve this, we verify whether the resulting graph respects Bhattacharya et al. (2021)’s algebraic characterization of the space of AGs. More specifically, any AG $\mathcal{G}$ is characterized by an adjacency matrix $\mathbf{A}_{d}\in\mathbb{R}^{n\times n}$ for its directed edges and another adjacency matrix $\mathbf{A}_{b}\in\mathbb{R}^{n\times n}$ for its bidirected edges, adhering to:
$\text{trace}(\exp\{\mathbf{A}_{d}\})-n+\mathbf{1}^{T}(\exp\{\mathbf{A}_{d}\}\odot\mathbf{A}_{b})\mathbf{1}=0,$ (8)
in which $\mathbf{1}$ is a $n$-dimensional unit vector and $\odot$ denotes the Hadamard (elementwise) product of matrices.
Score-based belief. We propose using a strictly decreasing transformation of a score function $U$ as the reward $R$ for AGFN. More precisely, we define $R$ as
$R(\mathcal{G})=\exp\left\{\frac{\mu-U(\mathcal{G})}{\sigma}\right\}$ (9)
for given constants $\mu\in\mathbb{R}$ and $\sigma\in\mathbb{R}^{+}$ that ensure numerical stability (Zhang et al., 2023). In practice, we sample $S$ AGs $\{\mathcal{G}^{(s)}\}_{s=1}^{S}$ from an untrained AGFN, and set $\mu=\nicefrac{{1}}{{S}}\sum_{s}U(\mathcal{G}^{(s)})$ and $\sigma=\sqrt{\nicefrac{{1}}{{S}}\sum_{s}(U(\mathcal{G}^{(s)})-\mu)^{2}}$.
Score for linear Gaussian models. Since we focus on linear Gaussian models, we choose the extended Bayesian Information Criterion Foygel and Drton (2010) as our score function. Specifically, for any AG $\mathcal{G}=(\mathbf{V},\mathbf{E})$:
$U(\mathcal{G})$ $=-2l_{N}(\hat{\mathbf{B}},\hat{\boldsymbol{\Omega}})+|\mathbf{E}|\log n+2|\mathbf{E}|\log|\mathbf{V}|,$ (10)
in which $(\hat{\mathbf{B}},\hat{\boldsymbol{\Omega}})$ is the MLE estimate of model parameters (see eq. 3) obtained using the residual iterative conditional fitting algorithm Drton et al. (2009).
Forward flow. We use a Graph Isomorphism Network (GIN) Xu et al. (2019) $\Phi$ to compute a $d$-dimensional representation for each node in the AG $\mathcal{G}_{t}$ at the $t$-th step of the generative process and use sum pooling to get an embedding for $\mathcal{G}_{t}$. Then, considering $\mathcal{A}_{t}$ as the space of feasible actions at $\mathcal{G}_{t}$ (i.e., those leading to an AG), we use an MLP $\phi\colon\mathbb{R}^{d}\rightarrow\mathbb{R}^{|\mathcal{A}_{t}|}$ with a softmax activation at its last layer to map $\mathcal{G}_{t}$’s embedding to a distribution over $\mathcal{A}_{t}$. More precisely, given $\mathbf{H}^{(t)}=\Phi(\mathcal{G}_{t})\in\mathbb{R}^{|\mathbf{V}|\times d}$, we compute
$\mathbf{p}=\phi\left(\sum_{v\in\mathbf{V}}\mathbf{H}_{v}^{(t)}\right)$ (11)
as the probability distribution over the feasible actions at $\mathcal{G}_{t}$.
Backward flow. Backward actions correspond to removing edges. Following Shen et al. (2023), we parametrize the backward flow $F_{B}$ with an MLP and alternate between updating $\pi_{F}$ and $\pi_{F_{B}}$, using gradient-based optimization.
4 Human-in-the-Loop Causal Discovery
Concluding the training, we propose leveraging the AGFN-generated samples to design questions for the expert that optimize the reduction of entropy in the distribution $p_{\theta}(\mathcal{G})$ over the space of AGs. Then, we use the human feedback to update $p_{\theta}(\mathcal{G})$, and iteratively repeat the process. The following paragraphs describe i) how we model human feedback, ii) how we update our belief over AGs given human responses, and iii) our experimental design strategy for expert inquiry.
Modeling human feedback.
We assume humans are capable of answering questions regarding the ancestral relationship between pairs of random variables. In this case, we model their prior knowledge on a relation $r=\{U,V\}$ between nodes $U,V\in\mathbf{V}$ as a categorical distribution over a random variable denoted $\omega_{r}$. Fix an arbitrary total order $<$ in $\mathbf{V}$. By definition, $\omega_{r}=1$ if there is no edge between $U$ and $V$; $\omega_{r}=2$ if $U$ is ancestor of $V<U$; $\omega_{r}=3$ if $V$ is ancestor of $U>V$; and $\omega_{r}=4$ if there is a bidirected edge between $U$ and $V$. Since the human has access to our AGFN before being probed for the first time, we set $\rho_{r,k}=p_{\theta}(\omega_{r}=k)$ as the prior probability of $\omega_{r}=k$. Moreover, we consider that the expert’s feedback $f_{r}\in\{1,2,3,4\}$ on the relation $r$ is a noisy realization of the true, unobserved value of the relation feature $\omega_{r}$ under the expert’s model. Putting all elements together results in a two-level Bayesian hierarchical scheme for categorical data:
$\omega_{r}$ $\sim\text{Cat}(\boldsymbol{\rho}_{r}),$ (12)
$f_{r}|\omega_{r}$ $\sim\text{Cat}\left(\delta_{\omega_{r}}\cdot\pi+(\mathbf{1}-\delta_{\omega_{r}})\cdot\left(\frac{1-\pi}{3}\right)\right),$ (13)
in which $\boldsymbol{\rho}_{r}=(\rho_{r,1},\rho_{r,2},\rho_{r,3},\rho_{r,4})$ represents our prior beliefs about the relations’ features, $\pi\in[0,1]$ reflects the reliability of the expert’s feedback, and $\delta_{k}$ is the $k$-th canonical basis of $\mathbb{R}^{4}$. Conveniently, the posterior distribution of the relation feature $\omega_{r}$ given the feedback $f_{r}$ is a categorical distribution parametrized by
$\frac{\boldsymbol{\rho}_{r}}{\eta_{r}}\odot\left(\pi\cdot\delta_{f_{r}}+\left(\frac{1-\pi}{3}\right)\cdot(\mathbf{1}-\delta_{f_{r}})\right),$ (14)
with $\eta_{r}=\rho_{r,f_{r}}\cdot\pi+\left(\frac{1-\pi}{3}\right)\cdot(1-\rho_{r,f_{r}})$.
Figure 3: Sampling quality. The reward-induced marginal distribution over graph features is adequately approximated by the marginal distribution learned by the GFlowNet.
Updating beliefs.
We update our AGFN by weighing it by our posterior over the expert’s knowledge, described in the previous paragraph, similarly to a product-of-experts approach Hinton (2002). For this, let $\mathbf{f}_{K}=(f_{r_{k}})_{1\leq k\leq K}$ be the sequence of $K$ feedbacks issued by the expert and define our novel belief distribution $q(\mathcal{G};\mathbf{f}_{K})$ over the space of AGs as
$q(\mathcal{G};\mathbf{f}_{K})\propto p_{\theta}(\mathcal{G})\prod_{1\leq k\leq K}p(\omega_{r_{k}}|f_{r_{k}}).$ (15)
Importantly, we use $p_{\theta}$ as proposal distribution in importance (re-)sampling Gordon et al. (1993) to approximately sample from $q(\mathcal{G};\mathbf{f}_{K})$ — or to approximate the expected value of a test function. More precisely, we estimate the value of a function $h$ over the space of AGs as:
$\mathbb{E}_{q}[h(\mathcal{G})]\approx\sum_{t=1}^{T}c^{-1}\frac{q(\mathcal{G}^{(t)};\mathbf{f}_{K})}{p_{\theta}(\mathcal{G}^{(t)})}h(\mathcal{G}^{(t)})$ (16)
with $(\mathcal{G}^{(t)})_{t=1}^{T}\sim p_{\theta}(\mathcal{G})$ and $c=\sum_{t=1}^{T}\nicefrac{{q(\mathcal{G}^{(t)};\mathbf{f}_{K})}}{{p_{\theta}(\mathcal{G}^{(t)})}}$.
Active knowledge elicitation.
To make the most out of possibly costly human interactions, we query the human about the relation that maximally reduces the expected cross-entropy between our belief over AGs before and after human feedback. More precisely, we define an acquisition function $a_{k}:{{\mathbf{V}}\choose{2}}\rightarrow\mathbb{R}$ for the $k>1$-th inquiry as:
$a_{k}(r)=-\mathbb{E}_{f_{r}\sim p(\cdot|\mathbf{f}_{K})}\big{[}\mathbf{H}\left(q(\mathcal{G};\mathbf{f}_{K},f_{r}),q(\mathcal{G};\mathbf{f}_{K})\right)\!\big{]}$ (17)
in which $p(f_{r}|\mathbf{f}_{K})$ is the posterior predictive distribution according to the user model, $q_{0}\propto R$ and $\mathbf{H}(\cdot,\cdot)$ is the cross-entropy. Then, we maximize this acquisition to select which relation $\tilde{r}_{k}$ we will probe the expert, i.e.:
$\tilde{r}_{k}=\operatorname*{arg\ max}_{r\in{{\mathbf{V}}\choose{2}}}a_{k}(r).$ (18)
As aforementioned, we use importance sampling with $q_{0}$ as a proposal to estimate the acquisition function $a_{k}$. This allows us to leverage AGFN samples and effectively avoid the need for retraining them. It is worth mentioning that because $\mathbf{H}(p,p^{\prime})\geq\mathbf{H}(p,p)$ for any two distributions $p$ and $p^{\prime}$ of the same support, our strategy is equivalent to minimizing an upper bound on the entropy of $q_{k}$. Also, different from acquisitions based on information gain or mutual information, approximating Equation 17 via Monte Carlo does not require exhaustive integration over the space of AGs to yield asymptotically unbiased estimates — see Appendix.
chain4 IV collfork
SHD BIC SHD BIC SHD BIC
FCI $3.03{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.13}}$ 1.13 $5481.33{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.69}}$ 2.69 $\mathbf{3.75}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 0.64}}$ 0.64 $\mathbf{5426.18}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.74}}$ 1.74 $6.26{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.20}}$ 1.20 $5433.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 6.94}}$ 6.94
GFCI $\mathbf{2.24}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 0.64}}$ 0.64 $\mathbf{5479.77}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.75}}$ 1.75 $4.21{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 0.96}}$ 0.96 $5427.09{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.85}}$ 2.85 $\mathbf{5.23}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.08}}$ 1.08 $\mathbf{5431.67}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 7.91}}$ 7.91
DCD $3.38{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.30}}$ 1.30 $5482.97{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 5.16}}$ 5.16 $5.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.23}}$ 1.23 $5429.51{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 4.37}}$ 4.37 $6.02{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.22}}$ 1.22 $5436.84{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 9.41}}$ 9.41
[…]
AGFN (ours) $\mathbf{6.04}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.12}}$ 2.12 $\mathbf{5494.67}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 37.08}}$ 37.08 $8.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.04}}$ 2.04 $\mathbf{5456.16}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 52.25}}$ 52.25 $\mathbf{6.58}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.34}}$ 2.34 $\mathbf{5478.01}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 40.36}}$ 40.36
Table 1: Average SHD and BIC. The denotes methods yielding point estimates. We use Bootstrap to report the mean and average standard deviation for these. For N-ADMG and AGFN, we estimate the quantities using $100$k samples.
5 Experiments
Our experiments have three objectives. First, we validate that AGFN can accurately learn the target distribution over the space of AGs. Second, we show that AGFN performs competitively with alternative methods on three data sets. Third, we attest that our experimental design for incorporating the expert’s feedback efficiently reduces the uncertainty over AGFN’s distribution. We provide further experimental details in the Appendix. Code is available in the supplement.
5.1 Distributional Assessment of AGFN
Data. Since violations of faithfulness are more likely in dense graphs Uhler et al. (2012), we create $20$ $5$-node random graphs Uhler et al. (2012) from a directed configuration model (Newman, 2010) whose in- and out-degrees are uniformly sampled from $\{0,1,2,3,4\}$. We draw $500$ independent samples from a structure-compatible linear Gaussian SCM with random parameters for each graph.
[…]
Results. Figure 3 shows that the AGFN adequately approximates the theoretical distribution induced by the reward in Equation 9. Furthermore, AGFNs induce distributions over BIC and SHD values that closely resemble those induced by $p(\mathcal{G})\propto R(\mathcal{G})$. We also note an important improvement over the prior art on probabilistic CD (N-ADMG): we found that over $60\%$ of its samples were non-ancestral and that this method was of little use for making inferences over AGs. Meanwhile, AGFN does not sample non-ancestral graphs.
5.2 Comparison with SOTA CD algorithms
(a) chain4
(b) IV
(c) collfork
Figure 4: Ancestral graphs representing the data generating models for the three considered datasets in Table 1.
Data. We generate $10$ datasets with $500$ independent samples from the randomly parametrized linear Gaussian SCMs corresponding to the canonical causal diagram Richardson and Spirtes (2002) in each AG depicted in Figure 4. Unshielded colliders and discriminating paths are fundamental patterns in the detection of invariances by CD algorithms under latent confounding Spirtes and Richardson (1997); Zhang (2008b). Thus, we consider the following 4-node causal diagrams with increasingly difficult configurations: (i) chain4, a chain without latent confounders; (ii) collfork, a graph with triplets involving colliders and non-colliders under latent confounding, and (iii) IV, a structure with a discriminating path for $Z$: $W\rightarrow X\leftarrow Z\rightarrow Y$.
Baselines. We compare AGFN with four notable CD methods: FCI (Spirtes et al., 2001; Zhang, 2008b), GFCI (Ogarrio et al., 2016), DCD (Bhattacharya et al., 2021), and N-ADMG Ashman et al. (2023). The baselines span four broad classes of CD methods. FCI is a seminal constraint-based CD algorithm that learns a PAG consistent with conditional independencies entailed by statistical tests. GFCI is a hybrid CD algorithm that learns a PAG by first obtaining an approximate structure using FGS (Ramsey, 2015) (a BIC-score-based search algorithm for causally sufficient scenarios) and then by applying FCI to identify possible confounding and remove some edges added by FGS. DCD casts CD as continuous optimization with differentiable algebraic constraints defining the space of AGs and uses gradient-based algorithms to solve it. N-ADMG computes a variational approximation of the joint posterior distribution over the space of bow-free causal diagrams (Nowzohour et al., 2017) associated with non-linear SCMs with additive noise. While N-ADMG focuses on a more restricted setting compared to AGs, it offers some uncertainty quantification in the variational pohttps://www.overleaf.com/project/650454a084b798332af29ebesterior, making it more closely comparable to our approach. We rigorously follow the experimental guidelines in the original works.
Experimental setup We train AGFN on each dataset and use it to sample 100k graphs. We also apply FCI, GFCI, and DCD to $100$ bootstrapped resamplings of each dataset to emulate confidence distributions induced by these algorithms. To compare the algorithms’ outputs, we compute the sample mean and standard deviation of the BIC and SHD at the PAG level. Specifically, we compute the SHD between the ground-truth PAG and each estimated PAG obtained by each method. If the output is a PAG member (as for DCD, N-ADMG, and AGFN) we use FCI to transform the output using these graphs as oracles for conditional independencies. Furthermore, we directly compute the BIC for the outputs, as all PAG members are asymptotically score-equivalent.
Results. Table 1 compares AGFN against baseline CD algorithms. Notably, our method consistently outperforms the only probabilistic baseline in the literature (N-ADMG) in terms of both SHD and BIC. As expected, however, the average BIC and SHD induced by AGFN are larger than those induced by the bootstrapped versions of the non-probabilistic algorithms, and the variances are greater; this is due to the inherent sampling diversity of our method and the resulting generation of possibly implausible samples. Indeed, Table 2 shows that the three most rewarding samples from AGFN are as good as (and sometimes better than) the other CD algorithms. Results for N-ADMG comprise the three most frequent samples from the variational distribution.
chain4 IV collfork
FCI $2.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.00}}$ 2.00 $3.83{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.90}}$ 2.90 $5.43{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.87}}$ 1.87
GFCI $\mathbf{1.50}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.63}}$ 1.63 $3.63{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 3.16}}$ 3.16 $5.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.11}}$ 2.11
DCD $2.27{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.46}}$ 1.46 $4.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.17}}$ 2.17 $5.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\pm\scriptstyle 2.13}}$ 2.13
N-ADMG (top 3) $4.38{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 0.81}}$ 0.81 $6.08{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.77}}$ 1.77 $6.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 0.93}}$ 0.93
AGFN (top 3) $2.00{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.55}}$ 1.55 $\mathbf{3.50}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 3.29}}$ 3.29 $\mathbf{4.90}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.70}}$ 2.70
Table 2: SHD for point estimates. The mean SHD of the top-3 AGFN draws is comparable to or better than baselines.
Figure 5: CD with simulated human feedback. The top/bottom row shows the mean SHD/BIC of AGFN samples as a function of human interactions. Probing the expert about the edge that minimizes the mean cross-entropy leads to a faster decrease in BIC compared to a random strategy. The SHD decreases similarly in both cases. Results reflect the outcomes of 30 simulations.
5.3 Simulating humans in the loop
Data. We follow the procedure from Section 5.1 to generate graphs with $4$, $6$, $8$ and $10$ nodes. We draw $500$ samples from a compatible linear Gaussian SCM and use them to train an AGFN. Then, we follow our active elicitation strategy from Section 4 to probe simulated humans, adhering to the generative model described in the same section, with $\pi=0.9$.
Setup. Since we are the first to propose an optimal design for expert knowledge elicitation, there are no baselines to compare AGFN against. That being said, we aim to determine whether the inclusion of expert feedback enhances the concentration of the learned distribution around the true AG, and evaluate the effectiveness of our elicitation strategy. To do so, we measure SHD to the true AG and BIC as a function of the number of expert interactions.
Results. Figure 5 shows that incorporating expert feedback substantially decreases the expected SHD and BIC under our belief over AGs. On the one hand, the remarkable decrease in expected SHD shows that our belief becomes increasingly focused on the true AG as we iteratively request the expert’s feedback, regardless of the querying strategy. On the other hand, the second row shows that our querying strategy results in a substantial decrease in the BIC, demonstrating a faster reduction than random queries. This validates the notion that some edges are more informative than others, and we should prioritize them when probing the expert.
6 Related Work
CD under latent confounding. Following the seminal works by Spirtes et al. (2001) and Zhang (2008b) introducing the complete FCI, a variety of works have emerged. Among them are algorithms designed for sparse scenarios, including RFCI (Colombo et al., 2012) and others (Silva, 2013; Claassen et al., 2013). Notably, Silva (2013)’s framework uses a Bayesian approach to CD of Gaussian causal diagrams based on sparse covariance matrices. Nonetheless, it requires sampling one edge at a time and relies on numerical heuristics that might effectively alter the posterior we are sampling from. Colombo et al. (2012) introduced the conservative FCI to handle conflicts arising from statistical errors in scenarios with limited data, even though it yields less informative results. Subsequent efforts to improve reliability led to the emergence of constraint-based CD algorithms based in Boolean satisfiability (Hyttinen et al., 2014; Magliacane et al., 2016), although they are known to scale poorly on $|\mathbf{V}|$ (Lu et al., 2021). In another paradigm, score-based search algorithms rank MAGs according to goodness-of-fit measures, commonly using BIC for linear Gaussian SCMs (Triantafillou and Tsamardinos, 2016; Zhalama et al., 2017a; Rantanen et al., 2021). There are also hybrid approaches that combine constraint-based strategies to reduce the search space, such as GFCI (Ogarrio et al., 2016), M3HC (Tsirlis et al., 2018), BCCD (Claassen and Heskes, 2012), and GSPo (Bernstein et al., 2020). Continuous optimization approaches have recently emerged as a novel approach to score-based CD, such as DCD (Bhattacharya et al., 2021) and N-ADMG Ashman et al. (2023).
CD with expert knowledge. Previous works on CD have explored various forms of background knowledge. This includes knowledge on edge existence/non-existence Meek (1995b), ancestral constraints Chen et al. (2016), variable grouping Parviainen and Kaski (2017), partial order Andrews (2020) and typing of variables (Brouillard et al., 2022). Incorporating expert knowledge is pivotal to reducing the search space and the size of the learned equivalence class. However, due to significant challenges, up to date, there are only a few works trying to integrate human knowledge into CD within the context of latent confounding (Andrews, 2020; Wang et al., 2022). These works operate under the assumption of perfect expert feedback. In contrast, our contribution is novel in that it confronts the challenges of real-world situations where expert input might be inaccurate.
7 Discussion
We presented AGFN, the first probabilistic CD method that accounts for latent confounding and incorporates potentially noisy human feedback in the loop. AGFN samples AGs according to a score function, quantifying the uncertainty in the learning process. Furthermore, it can leverage human feedback in an optimal design strategy, efficiently reducing our uncertainty on the true data-generating model.
This work is focused on linear Gaussian models, using BIC as our score. However, the implementation of AGFNs is not restricted by this choice. In principle, we could replace the BIC with alternative score functions that are more appropriate for different types of variables, e.g. for discrete data (Drton and Richardson, 2008). It is also important to highlight that our framework does not require retraining the AGFN after we see human feedback. Moreover, AGFN is a GPU-powered algorithm, and while we used only one GPU in our experiments, it is possible to greatly accelerate AGFN by using cluster architectures with multiple GPUs.
By offering uncertainty-quantified CD together with a recipe for including humans in the loop, we expect AGFNs will significantly enhance the accuracy and reliability of CD, especially in real-world domains. Moreover, AGFNs bring a novel perspective to developing more comprehensive tools for downstream causal tasks Bareinboim and Pearl (2016), as the resulting distribution encodes knowledge from data and human feedback while accounting for epistemic uncertainty. For example, methods for causal reasoning that currently rely on a single AG Zhang (2008a); Jaber et al. (2022) could exploit this distribution to incorporate a richer understanding of uncertainty and knowledge, thereby enhancing their robustness and reliability.
Acknowledgments
Diego Mesquita acknowledges the support by the Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro FAPERJ (SEI-260003/000709/2023), the São Paulo Research Foundation FAPESP (2023/00815-6), the Conselho Nacional de Desenvolvimento Científico e Tecnológico CNPq (404336/2023-0), and the Silicon Valley Community Foundation through the University Blockchain Research Initiative (Grant #2022-199610). António Góis acknowledges the support by Samsung Electronics Co., Ldt. Adèle Ribeiro and Dominik Heider were supported by the LOEWE program of the State of Hesse (Germany) in the Diffusible Signals research cluster and by the German Federal Ministry of Education and Research (BMBF) [031L0267A] (Deep Insight). Samuel Kaski was supported by the Academy of Finland (Flagship programme: Finnish Center for Artificial Intelligence FCAI), EU Horizon 2020 (European Network of AI Excellence Centres ELISE, grant agreement 951847), UKRI Turing AI World-Leading Researcher Fellowship (EP/W002973/1). We also acknowledge the computational resources provided by the Aalto Science-IT Project from Computer Science IT.
Appendix A Additional Related Works
Generative Flow Networks (GFlowNets; Bengio et al., 2021a, b) are generative models that sample discrete composite objects from an unnormalized reward function. They have been successfully used to sample various structures such as protein sequences (Jain et al., 2022) and schedules (Zhang et al., 2023). They have also been used to train energy-based models (Zhang et al., 2022). In the field of structure learning, they have been applied to Bayesian networks — more specifically to sample a posterior over DAGs in linear Gaussian networks, although without accounting for unobserved confounding (Deleu et al., 2022). Recently, Deleu et al. (2023) proposed an extension to jointly infer the structure and parameters, also grounded in the assumption of causal sufficiency. It is worth highlighting that training GFlowNets in these scenarios presents optimization challenges, resulting in the utilization of a variety of loss functions (Shen et al., 2023). Moreover, Lahlou et al. (2023) proposed an extension of GFlowNets to continuous domains.
Appendix B Cross-entropy acquisition
The expected mutual information and the information gain are the most widely used information-theoretic measures to actively interact with a human and choose the most informative data points to be labeled Ryan et al. (2015). However, we instead use the negative expected cross-entropy between the current and updated beliefs as the acquisition function of our experimental design (see eq. 17). As we show next, the approximation of both the mutual information and the information gain is intrinsically dependent upon the estimation of the log-partition of the updated beliefs over the space of ancestral graphs. Doing so is computationally intensive, and we would either need to use a Monte Carlo estimator of the integrals or use some posterior approximation — in both cases, leading to asymptotically biased estimates of the acquisition. In contrast, we can easily leverage AGFN samples to compute asymptotically unbiased estimates of our acquisition function. The next paragraphs provide further details.
Mutual information.
The mutual information between two random variables $X$ and $Y$ with joint distribution $p(X,Y)$ and marginal distributions $p(X)$ and $p(Y)$ is
$I(X,Y)=\mathcal{D}_{KL}[p(X,Y)||p(X)\otimes p(Y)],$ (19)
in which $\mathcal{D}_{KL}$ is the Kullback-Leibler divergence. In this context, an alternative approach to our experimental design for active knowledge elicitation would consist in iteratively maximizing the expected mutual information between the observed samples, $\mathcal{G}$, and the elicited feedback, $f_{K}$, to select the relation about which the expert would provide feedback. More specifically, we could choose
$r_{K+1}=\operatorname*{arg\ max}_{r\in{\mathbf{V}\choose 2}}\mathbb{E}_{f_{r}\sim p(\cdot|\mathbf{f}_{K})}[I(\mathcal{G},f_{r})],$ (20)
in which
$I(\mathcal{G},f_{r})=\mathcal{D}_{KL}[q(\mathcal{G},f_{r}|\mathbf{f}_{K})||q(\mathcal{G}|\mathbf{f}_{K})\otimes p(f_{r}|\mathbf{f}_{K})],$ (21)
at each interaction with the expert. Nonetheless, note that
$q(\mathcal{G},f_{r}|\mathbf{f}_{K})=q(\mathcal{G}|\mathbf{f}_{K+1})p(f_{r}|\mathbf{f}_{K})$
$=c_{K+1}(f_{r})p_{\theta}(\mathcal{G})\left(\prod_{1\leq k\leq K+1}p(\omega_{r_{k}}|f_{r_{k}})\right)\cdot p(f_{r}|\mathbf{f}_{K}),$
with $f_{r_{K+1}}=f_{r}$ and
$c_{K+1}(f_{r})=\left(\!\!\sum_{\mathcal{G}}p_{\theta}(\mathcal{G})\!\!\left(\prod_{1\leq k\leq K+1}p(\omega_{r_{k}}|f_{r_{k}})\!\!\right)\!\!\right)^{-1}$ (22)
as the partition function of our updated beliefs. Note also that Equation 21 entails computing the entropy of $q(\mathcal{G},f_{r}|\mathbf{f}_{K})$. Thus, the selection criterion in eq. 20 requires an accurate estimate of $\log c_{K+1}(f_{r})$ — which is well-known for being a difficult problem Ma et al. (2013) — and the Monte Carlo estimator for the log-partition function is asymptotically biased.
Information gain.
The expected information gain of an elicitation is defined as the expected KL divergence between our updated and current beliefs over ancestral graphs. This approach is widely employed in Bayesian experimental design Ryan et al. (2015). In our framework, the information gain resulting from a feedback $f_{r}$ is
$\text{IG}_{K}(f_{r})=\mathcal{D}_{KL}[q(\mathcal{G}|\mathbf{f}_{K}\cup f_{r})||q(\mathcal{G}|\mathbf{f}_{K})],$ (23)
which yields the criterion
$r_{K+1}=\operatorname*{arg\ max}_{r\in{\mathbf{V}\choose 2}}\mathbb{E}_{f_{r}\sim p(\cdot|\mathbf{f}_{K})}\left[\text{IG}_{K}(f_{r})\right].$ (24)
Nonetheless, eq. 24 suffers from the same problems of eq. 20: it requires approximating the logarithm of the partition function $c_{K+1}(f_{r})$ of a distribution over the combinatorially large space of ancestral graphs, which is notably very challenging to estimate. Indeed, as
$\mathcal{D}_{KL}[q(\mathcal{G}|\mathbf{f}_{K+1})||q(\mathcal{G}|\mathbf{f}_{K})]=\underset{\mathcal{G}\sim q(\cdot|\mathbf{f}_{K+1})}{\mathbb{E}}\left[\log\frac{q(\mathcal{G}|\mathbf{f}_{K+1})}{q(\mathcal{G}|\mathbf{f}_{K})}\right]$
$=\underset{\mathcal{G}\sim q(\cdot|\mathbf{f}_{K+1})}{\mathbb{E}}\left[\log p(f_{r}|\omega_{r})+\log c_{K+1}(f_{r})-\log c_{K}\right],$
with $f_{r_{K+1}}=f_{r}$, $c_{K}$ as the partition function of $q(\cdot|\mathbf{f}_{K})$ — that does not depend upon $f_{r}$ —, and $c_{K+1}(f_{r})$ defined in eq. 22, the estimation of the information gain is inherently dependent upon the estimation of the log-partition function.
Cross-entropy.
The cross-entropy between our updated and current beliefs is an intuitively plausible and practically useful strategy to interact with an expert efficiently. In fact, since
$\begin{split}\mathbf{H}&[q(\cdot|\mathbf{f}_{K+1}),q(\cdot|\mathbf{f}_{K})]\\ &=\underset{\mathcal{G}\sim q(\cdot|\mathbf{f}_{K+1})}{\mathbb{E}}[-\log q(\mathcal{G}|\mathbf{f}_{K})]\\ &=\underset{\mathcal{G}\sim q(\cdot|\mathbf{f}_{K+1})}{\mathbb{E}}\left[-\log p_{\theta}(\mathcal{G})-\sum_{1\leq k\leq K}\log p(\omega_{r_{k}}|f_{r_{k}})-l_{K}\right]\end{split},$
in which $l_{K}$ is the log-partition function of the distribution $q(\cdot|\mathbf{f}_{K})$. Further, the cross-entropy depends exclusively upon i) the logarithm of the samples’ rewards, $\log p_{\theta}(\mathcal{G})$, which is readily computed within AGFN’s generative process, and ii) the posterior distribution over the relations’ features $\omega_{r}$ given the expert’s feedbacks $f_{r}$, which is available in closed form. Hence, the previously mentioned expectation is unbiasedly and consistently estimated by our importance sampling scheme. Furthermore, our empirical findings in fig. 5 suggest that the cross-entropy yields good results and consistently outperforms a uniformly random strategy with respect to the BIC score.
Appendix C Experimental details
We lay out the experimental and implementational details of our empirical analysis in the next subsections. In Section C.1, we describe the specific configurations of the CD algorithms that we compared with our method in table 2. Then, we consider in section C.2 some practical guidelines and architectural specifications that enable us to train and make inferences with AGFN efficiently. Finally, we contemplate in section C.3 the algorithmic details for simulating the expert’s feedback according to our model for active knowledge elicitation.
C.1 Baselines
FCI.
For the results in table 1, we first estimated a PAG using the stable version of FCI, which produces a fully order-independent final skeleton (Colombo et al., 2014). To identify conditional independencies, we used Fisher’s Z partial correlation test with a significance level of $\alpha=0.05$. The BIC score associated with the PAG estimated by the FCI was computed as the BIC of a randomly selected maximal AG (MAG) within the equivalence class characterized by such PAG. The maximality of an AG depends on the absence of inducing paths between non-adjacent variables, which are paths where every node along it (except the endpoints) is a collider and every collider is an ancestor of an endpoint (Rantanen et al., 2021). This ensures that in the MAG every non-adjacent pair of nodes is m-separated by some set of other variables. Importantly, Markov equivalent MAGs exhibit asymptotic equivalence in terms of BIC scores (Richardson and Spirtes, 2002). As a result, the choice of a random MAG does not disrupt the validity of our results.
GFCI.
Similarly, we applied GFCI with an initial search algorithm (FGS) based on the BIC score and the subsequent application of the FCI with conditional independencies identified by the Fisher’s Z partial correlation test with a significance level $\alpha=0.05$. This was performed for all datasets listed in table 1. Also similar to the procedure adopted with the FCI, the BIC score associated with the estimated PAG was computed as the BIC of a randomly selected MAG within the equivalence class characterized by such PAG.
DCD.
We adhered to the instructions provided in the official repository [^1] Available online at https://gitlab.com/rbhatta8/dcd. to apply the DCD method on the datasets in table 1. The SHD was obtained between the ground-truth PAG and the PAG corresponding to the estimated ADMG (i.e., the one obtained via FCI by using the d-separations entailed by the estimated ADMG as an oracle for conditional independencies). On the other hand, the BIC was computed for the estimated ADMG directly.
N-ADMG.
To estimate the parameters of the variational distribution defined by N-ADMG, we executed the code provided at the official repository [^2] Available online at https://github.com/microsoft/causica/releases/tag/v0.0.0. For fairness, we used the same hyperparameters and architectures reported in their original work Li et al. (2023); in particular, we trained the models for $30$k epochs. After this, we sampled $100$k graphs from the learned distribution. It is worth mentioning that the constraints of bow-free ADMG are guaranteed in the N-ADMG samples only in an asymptotic sense. Thus, we manually removed any cyclic graphs from the learned distribution. Then, we proceeded exactly as with DCD to estimate both the average SHD and the average BIC under the variational distribution.
C.2 Implementational details for AGFN
Figure 6: Tempered rewards. Training AGFN to sample from increasingly cold distributions (eq. 25) enables us to increase the proportion of high-scoring graphs (i.e., with a low BIC-score) with the drawback of reducing the AGFN’s sampling diversity.
Masking.
To ensure AGFN only samples ancestral graphs, we keep track of a binary mask $\mathbf{m}_{t}$ that indicates which actions lead to a valid state at the iteration $t$ of the generative process; this mask defines the support of the policy evaluated at the corresponding state. In more detail, let $\mathbf{y}_{t}$ be the last layer embedding (prior to a softmax) at iteration $t$ of the neural network used to parametrize the forward flow of AGFN. The probability distribution over the space of feasible actions is then
$\mathbf{p}_{t}=\text{Softmax}\left(\mathbf{y}_{t}\odot\mathbf{m}_{t}+\epsilon\cdot(1-\mathbf{m}_{t})\right)$
for a large and negative constant $\epsilon$. We empirically verified that $\epsilon=-10^{5}$ is sufficient to avoid the sampling of non-ancestral graphs.
Exploratory policy.
During training, we must use an exploratory policy that (i) enables the exploration of yet unvisited states within the pointed DAG and (ii) exploits highly valuable and already visited states. To alleviate this phenomenon, we also draw trajectories from a uniform policy, which is a widespread practice in the literature (Bengio et al., 2021a; Deleu et al., 2022; Shen et al., 2023). More precisely, let $\text{Ch}(\mathcal{G}_{t})$ be the set of states (i.e., ancestral graphs) directly reachable from $\mathcal{G}_{t}$ and $\alpha\in[0,1]$. At each iteration $t$ of the generative process, we sample an action (either an edge to be appended to the graph or a signal to stop the process)
$a_{t}\sim(1-\alpha)\cdot\mathcal{U}(\text{Ch}(\mathcal{G}_{t}))+\alpha\cdot\pi_{F}(\cdot|\mathcal{G}_{t})$
and modify $\mathcal{G}_{t}$ accordingly. The parameter $\alpha$ quantifies the mean proportion of on-policy actions and represents a trade-off between choosing actions that lead to highly valuable states ($\alpha=1$) and actions that lead to unvisited states ($\alpha=0$). We fix $\alpha=\frac{1}{2}$ throughout the experiments. During inference, we set $\alpha=1$ to sample actions exclusively from the GFlowNet’s learned policy.
Detection of invalid states.
We use the algebraic condition in eq. 8 to check whether a graph $\mathcal{G}$ is ancestral. At each iteration of the generative process, we draw an action from the current exploratory policy and test the ancestrality of the updated graph; if it is not ancestral, we revert the sampled action and mask it. Importantly, this protocol guarantees that all graphs sampled from AGFN are ancestral.
Batch sampling.
We exploit batch sampling to fully leverage the power of GPU-based computing in AGFN. As both the maximum-log-likelihood-based reward and the validation of the states are parallelizable operations, we are able to distribute them across multiple processing units and efficiently draw samples from the learned distribution. Crucially, this end-to-end parallelization substantially improves the computational feasibility of our algorithm and is a notable feature generally unavailable in prior works Zhang (2008b); Ogarrio et al. (2016); Rantanen et al. (2021). We use a batch size of 256 for all the experiments — independently of the graph size.
Figure 7: Sensitivity of our active knowledge elicitation framework to the reliability of the expert. Each column represents either the expected SHD (top) or expected BIC (bottom) as a function of the degree of confidence $\pi\in[0,1]$ in the expert as a function of the number of feedbacks. As expected, the improvements entailed by the expert’s feedback become increasingly effective as we increase the expert’s reliability from $0.1$ to $0.9$. Results reflect the outcome of 30 scenarios simulated accordingly to algorithm 1 with a random canonical diagram $\mathcal{G}^{\star}$ with $5$ nodes. We used our active knowledge elicitation scheme to select the query at each iteration.
Training hyperparameters.
For AGFN’s forward flow, we use a Graph Isomorphism Network (GIN, Xu et al. (2019)) with $2$ layers to compute embeddings of dimension $256$. Then, we project these embeddings to a probability distribution using a three-layer MLP having leaky RELUs with a negative slope of $-0.01$ as activation functions. Correspondingly, we use an equally configured three-layer MLP to parametrize AGFN’s backward flow. For training, we use the Adam method for the stochastic optimization problem defined by the minimization of the loss in eq. 7. Moreover, we trained the neural networks for $3000$ epochs for the human-in-the-loop simulations (in which we considered graphs having up to $10$ nodes) and for $500$ epochs for both the assessment of the distributional quality of AGFN and the comparison of AGFN with alternative CD approaches.
Computational settings.
We trained the AGFNs for the experiments in fig. 3 and table 1 and fig. 5 for $500$ epochs in computers equipped with NVIDIA’s V100 GPUs. All the experiments were executed in a cluster of NVIDIA’s V100 GPUs and the algorithms were implemented using the machine learning framework PyTorch. To estimate the PAG corresponding to AGFN’s samples and compute the SHDs reported in table 1, we used the FCI’s implementation of the pcalg package in R considering the d-separations entailed by these samples as a criterion for conditional dependence.
C.3 Human in the loop
Algorithm 1 Simulating humans in the loop
$\{\mathcal{G}_{t}\}_{1\leq t\leq T}\text{ samples from AGFN}$, $\mathcal{G}^{*}=(\mathbf{V},E)$ true ancestral graph, $\pi$ reliability of the expert’s feedback
$p(\omega_{r}=k)\leftarrow\frac{1}{T}\sum_{1\leq t\leq T}1_{\{\omega_{r}=k\text{ in }G_{t}\}}\forall k\in[4],r\in{\mathbf{V}\choose 2}$
$\mathbf{f}\leftarrow\{\}$triangleright Set of feedbacks (answers)
$\boldsymbol{r}\leftarrow\{\}$triangleright Set of queries (questions)
$K\leftarrow 1$
$\omega_{r}^{*}\leftarrow\text{ relation }r\text{'s feature in }\mathcal{G}^{*}\ \forall r\in{\mathbf{V}\choose 2}$
while $\boldsymbol{r}\neq{\mathbf{V}\choose 2}$ do $\triangleright$ Iteratively request the feedback
     $r_{K}\leftarrow\underset{r\in{\mathbf{V}\choose 2}\setminus\boldsymbol{r}}{\operatorname*{arg\ max}}\ \underset{{f_{r}\sim p(\cdot)}}{\mathbb{E}}[-\mathbf{H}\left(q(\mathcal{G};\mathbf{f}\cup\{f_{r}\}),q(\mathcal{G};\mathbf{f})\right)]$
     $\mathbf{r}\leftarrow\mathbf{r}\cup\{r_{K}\}$
     $f_{K}\sim\text{Cat}\left(\pi\cdot\delta_{\omega_{r_{K}}^{*}}+\left(\frac{1-\pi}{3}\right)\cdot(1-\delta_{\omega_{r_{K}}^{*}})\right)$
     $\mathbf{f}\leftarrow\mathbf{f}\cup\{f_{K}\}$
     $K\leftarrow K+1$
end while
Algorithmic details.
We describe in algorithm 1 our procedure for simulating interactions with an expert. Initially, we estimate the marginal probabilities $p(\omega_{r}=k)$ of a relation $r$ displaying the feature $k\in\{1,2,3,4\}$ under AGFN’s learned distribution. This is our prior distribution. In algorithm 1, we denote $\{1,2,3,4\}$ by $[4]$. Then, we iteratively select the relation that maximizes our acquisition function; the simulated human thus returns a feedback that equals the selected relation’s true feature with probability $\pi$ or is otherwise uniformly distributed among the incorrect alternatives. Importantly, this iterative mechanism can be interrupted at any iteration and the collected feedbacks can be used to compute the importance weights necessary for estimating expectations of functionals under our updated beliefs.
Appendix D Additional experiments
Figure 8: Architectural design of AGFN. Top: The inductively biased parametrization of AGFN’s forward flow — based upon a GNN — enables the substantial reduction of the number of epochs required for training. Bottom: The use of a parametrized backward policy similarly enhances the training efficiency compared to a uniform policy. For both experiments, we considered $\mathcal{L}(\theta)<0.1$ as the early stopping criterion to interrupt AGFN’s training.
Trade-off between diversity and optimality in AGFN.
We may use tempered rewards to increase the frequency of high-scoring samples and thereby reduce the diversity of AGFN’s distribution. More precisely, we choose a temperature $T$ and consider
$R_{T}(\mathcal{G})=R(\mathcal{G})^{1/T}=\exp\left\{\frac{\mu-U(\mathcal{G})}{T\sigma}\right\}$ (25)
as the reward upon which the GFlowNet is trained; if $T\rightarrow 0$, the distribution $p_{T}\propto R_{T}$ converges to a point mass in $R(\mathcal{G})$’s mode and, if $T\rightarrow\infty$, $p_{T}$ converges to an uniform distribution. This approach resembles the simulated tempering scheme commonly exploited in Monte Carlo methods Marinari and Parisi (1992) and was previously considered in the context of GFlowNets by Zhang et al. (2023). Figure 6 shows that progressively cold distributions (i.e., with $T\rightarrow 0$) lead to progressively concentrated and decreasingly diverse samples. Notably, the use of cold distributions may be adequate if we are highly confident in our score and are mostly interested in high-scoring samples (e.g., as in Rantanen et al., 2021).
Figure 9: Human-aided AGFN significantly outperforms alternative CD algorithms. Updating AGFN’s distribution according to the feedback of an oracle substantially improves AGFN’s capacity to correctly identify the true ancestral graph; indeed, a single feedback is sufficient to yield results better than (or indistinguishable from) alternative CD algorithms. We select the sampled AG with the highest posterior reward as a point estimate of AGFN and use the same datasets listed in table 2. The plots summarize the results of $30$ HITL simulations using $\pi=0.9$ and an oracle as an expert (see algorithm 1).
Sensitivity analysis for different noise levels.
Figure 7 displays the effect of the feedback of an increasingly reliable expert over the expectations of both the SHD and the BIC. Notably, the usefulness of these feedbacks increases as the feedback noise decreases. This is expected as, for example, a completely unreliable expert consistently rules out only one of four possibilities for the features of each relation; then, there remains a great ambiguity, albeit not as much as there was prior to their feedback, about the true nature of the elicited causal relation. Moreover, this experiment highlights the potential to adjust the reliability parameter $\pi$ to incorporate knowledge into AGFN’s learned distribution regarding the non-existence of a particular relation, rather than its existence. More specifically, assume that the expert is certain that there is no directed edge from the variable $U$ to the variable $V$ in the underlying ancestral graph; for instance, a doctor may be certain that cancer ($U$) is not an ancestor (cause) of smoking ($V$), but may be uncertain about the definite relation between $U$ and $V$ (i.e., smoking may or may not cause cancer). To incorporate such knowledge into our model, one approach is to set a necessarily small reliability parameter $\pi$ (possibly, $\pi=0$) along with the improbable relation $U\rightarrow V$. This feedback will then be modeled as a relation unlikely to exist in the true ancestral graph. We emphasize that our model for the expert’s responses is straightforwardly extensible to accommodate multiple feedbacks about the same causal relation under different reliability levels.
Ablation studies.
Figure 8 shows the increase in the training efficiency due to our architectural designs for parametrizing both the forward and backward flows of AGFN. Noticeably, the use of a two-layer graph isomorphism network (GIN; Xu et al., 2019) with a 256-dimensional embedding for the forward flow entailed a decrease of more than 10x in the number of epochs required for successfully training AGFN; this highlights the effectiveness of an inductively biased architectural design for the parametrization of GFlowNet’s flows. Correlatively, the use of a parametrized backward flow significantly enhances the training efficiency of AGFN and emphasizes the inadequacy of a uniformly distributed backward policy pointed out in a previous work Shen et al. (2023).
Human-aided AGFN versus alternative CD methods.
Figure 9 exposes the significant enhancement of AGFN’s point estimates entailed by our HITL framework for CD. This underlines the usefulness of the elicited knowledge, which is simply incorporated into our model through a re-weighting of the reward function, enabling the identification of the true ancestral graph. In contrast, most alternative CD algorithms cannot be as easily adapted to include various forms of expert knowledge — and such incorporation, when it is possible, usually precedes any inferential process Andrews (2020) or assumes the knowledge is perfect Wang et al. (2022).

### v3

@@
-Human-in-the-Loop Causal Discovery under Latent Confounding using Ancestral GFlowNets
+Expert-Aided Causal Discovery of Ancestral Graphs
 Abstract
-Structure learning is the crux of causal inference.
-Notably, causal discovery (CD) algorithms are brittle when data is scarce, possibly inferring imprecise causal relations that contradict expert knowledge — especially when considering latent confounders.
-To aggravate the issue, most CD methods do not provide uncertainty estimates, making it hard for users to interpret results and improve the inference process.
-Surprisingly, while CD is a human-centered affair, no works have focused on building methods that both 1) output uncertainty estimates that can be verified by experts and 2) interact with those experts to iteratively refine CD.
-To solve these issues, we start by proposing to sample (causal) ancestral graphs proportionally to a belief distribution based on a score function, such as the Bayesian information criterion (BIC), using generative flow networks.
-Then, we leverage the diversity in candidate graphs and introduce an optimal experimental design to iteratively probe the expert about the relations among variables, effectively reducing the uncertainty of our belief over ancestral graphs.
-Finally, we update our samples to incorporate human feedback via importance sampling.
-Importantly, our method does not require causal sufficiency (i.e., unobserved confounders may exist).
-Experiments with synthetic observational data show that our method can accurately sample from distributions over ancestral graphs and that we can greatly improve inference quality with human aid.
+Causal discovery (CD) algorithms are notably brittle when data is scarce, inferring unreliable causal relations that may contradict expert knowledge, especially when considering latent confounders.
+Furthermore, the lack of uncertainty quantification in most CD methods hinders users from diagnosing and refining results.
+To address these issues, we introduce Ancestral GFlowNets (AGFNs).
+AGFN samples ancestral graphs (AGs) proportionally to a score-based belief distribution representing our epistemic uncertainty over the causal relationships.
+Building upon this distribution, we propose an elicitation framework for expert-driven assessment.
+This framework comprises an optimal experimental design to probe the expert and a scheme to incorporate the obtained feedback into AGFN.
+Our experiments show that: i) AGFN is competitive against other methods that address latent confounding on both synthetic and real-world datasets; and ii) our design for incorporating feedback from a (simulated) human expert or a Large Language Model (LLM) improves inference quality.
 1 INTRODUCTION
-Drawing conclusions about cause-and-effect relationships presents a fundamental challenge in various scientific fields and significantly impacts decision-making across diverse domains Pearl (2000).
-The importance of having structural knowledge, often encoded as a causal diagram, for conducting causal inferences is widely recognized, a concept made prominent by Cartwright (1989)’s dictum: “no causes in, no causes out”.
-When there is no objective knowledge to fully specify a causal diagram, causal discovery (CD) tools are instrumental in partially uncovering causal relationships among variables from, for example, observational data.
-Formally, let $\mathbf{V}=\{V_{1},V_{2},\dots,V_{n}\}$ be a set of $n$ observed variables and $\mathcal{D}$ be a dataset containing $|\mathcal{D}|=m$ samples for each $V_{i}\in\mathbf{V}$.
-A CD algorithm takes $\mathcal{D}$ as input and typically returns a single graph $\mathcal{G}=(\mathbf{V},\mathbf{E})$ with well-defined causal semantics, in which each node in $\mathbf{V}$ represents a variable $V_{i}\in\mathbf{V}$ and each edge in $\mathbf{E}$ encodes the possible underlying (causal/confounding) mechanisms compatible with $\mathcal{D}$.
-Figure 1: Human-in-the-loop probabilistic CD.
-We first train an AGFN to fit a data-informed belief over AGs.
-Then, we iteratively refine it by 1) questioning (Q) experts on the relation between a highly informative pair of variables and 2) updating the belief given the potentially noisy answers (A).
-The histograms on top of the edges show marginals over edge types (green denotes ground truth).
+Causal discovery (CD) methods are widespread in science as tools to uncover complex cause-and-effect relationships.
+These algorithms typically leverage observational data to infer a graphical representation of the class of models that are equally likely to have generated the data, known as the Markov Equivalence Class (MEC).
+However, they are known to be unreliable when data are scarce, as statistical relationships drawn from such data may not align with the true causal model.
+This mismatch defines what we call a violation of the faithfulness assumption (Zhang and Spirtes,, 2016).
+For instance, CD algorithms based on conditional independence tests may fall victim to false independence relations inferred due to a lack of statistical power.
+These statistical errors may propagate and trigger a chain reaction of erroneous edge orientations (Zhang and Spirtes’,, 2008; Zhalama et al., 2017b, ; Ng et al.,, 2021).
+Meanwhile, algorithms that maximize goodness-of-fit scores may infer a structure that is optimal for the observed data, but does not align with the ground-truth MEC (Ogarrio et al.,, 2016).
+In the presence of latent confounding, CD focuses on learning Ancestral Graphs (AGs), which conveniently encode causal structures while abstracting away unobserved confounders (Richardson and Spirtes,, 2002).
+Notably, learning AGs is particularly challenging as latent confounders both enlarge the space of models consistent with the data and weaken observable dependencies, thereby exacerbating model ambiguity and increasing the likelihood of faithfulness violations.
+CD algorithms such as FCI (Zhang, 2008b, ), GFCI (Ogarrio et al.,, 2016), ACI (Magliacane et al.,, 2016), and DCD (Bhattacharya et al.,, 2021) address latent confounding but lack uncertainty quantification, which is crucial under limited samples and possible faithfulness violations.
+Some extensions incorporate expert knowledge (Ribeiro et al.,, 2024; Wang et al.,, 2023; Andrews,, 2020; Ankan and Textor,, 2025), yet they require deterministic, noise-free inputs and support only a narrow range of knowledge types.
+A complete and seamless integration of probabilistic or noisy expert information remains absent.
+From a practitioner’s perspective, two questions are particularly relevant when interpreting CD results:
+  1. 1.
+    How much should results be trusted?
+  2. 2.
+    How can the quality of inference be improved?
+Two key components to address these questions are: 1) integrating uncertainty quantification into CD to enhance trustworthiness and transparency; and 2) developing mechanisms to incorporate experts in the loop.
+With this in mind, we propose a two-stage fully probabilistic CD framework called Ancestral GFlowNets (AGFNs).
+1.
+First, an AGFN approximates an energy-based distribution over AGs based on a given score function such as the Bayesian Information Criterion (Schwarz,, 1978).
+2.
+Then, it probes an uncertain expert (possibly an LLM) on the relationship between selected pairs of variables and uses their feedback for self-refinement.
+In doing so, AGFN can provide accurate responses even in the presence of unfaithful data, in contrast with standard CD algorithms.
+While most probabilistic CD rely on Markov Chain Monte Carlo (MCMC) methods (e.g., Silva and Ghahramanir,, 2009; Silva,, 2013), AGFNs sample AGs using Generative Flow Networks (GFlowNets; Bengio et al., 2021a, ; Bengio et al., 2021b, ), which are generative models designed to sample diverse modes without the mixing time limitations of MCMC and without relying on handcrafted proposals nor accept-reject steps (Bengio et al., 2021a, ).
+Importantly, the score function is a hyperparameter of AGFN, allowing users to select whichever function is most suitable for their data without requiring changes in our method.
+This extends the applicability of AGFN beyond linear Gaussian data.
+Additionally, although GFlowNets have been previously used for CD (Deleu et al.,, 2022, 2023), AGFNs are the first method that simultaneously account for latent confoundedness of the data and allow for the seamless integration of expert knowledge into the discovery process.
+For a detailed discussion on other related works, we refer the reader to Appendix E.
+To achieve this, we design an active elicitation strategy for AGFN for querying an expert about the existence and nature (whether ancestral or confounding) of an edge between the most informative pair of variables (Ryan et al.,, 2015).
+Subsequently, we update AGFN’s beliefs based on this (potentially noisy) feedback through a procedure that bypasses the need for model retraining.
+Figure 1 illustrates our iterative framework for expert-in-the-loop (EITL) probabilistic CD.
+Figure 1: EITL probabilistic CD.
+We iteratively refine the trained AGFN by i) querying (Q) experts on the relationship of a highly informative pair of variables and ii) updating beliefs based on their (noisy) answers (A).
+Histograms show marginals over edge types (green denotes ground truth).
 Notably, our belief increasingly concentrates on the true AG, $1\rightarrow 2\leftrightarrow 3$.
-This work focuses on recovering the structure of the underlying causal diagram when unobserved confounders may be at play.
-We propose to address this task by not only leveraging observational data but also by accounting for potentially noisy pieces of expert knowledge, otherwise unavailable as data.
-Throughout this work, we consider ancestral graphs (AGs) as surrogates for causal diagrams.
-AGs are particularly convenient since they encode latent confounding without explicitly invoking unobserved variables.
-Moreover, AGs capture all conditional independencies and ancestral relations among observed variables $\mathbf{V}$, as entailed by a causal diagram (Richardson and Spirtes, 2002).
-In the realm of CD from solely observational data, algorithms aim to construct a compact representation of the joint observational distribution $P(\mathbf{V})$, which implies a factorization as a product of conditional probabilities.
-Notably, multiple models may entail the same conditional independencies; in such cases, they are denoted as Markov-equivalent.
-As a result, these algorithms can only reconstruct the class of Markov-equivalent models (AGs), denoted as the Markov Equivalence Class (MEC) and typically represented by a Partial Ancestral Graph (PAG).
-Importantly, CD beyond the MEC by leveraging domain knowledge presents a critical challenge.
-Notably, there is no proper characterization of an equivalence class that accounts for knowledge stemming from both humans and data (Wang et al., 2022).
-There is a variety of algorithms for CD from observational data, primarily categorized into constraint- and score-based methods.
-The former uses (in)dependence constraints derived via conditional independence tests to directly construct a PAG representing the MEC.
-The latter uses a goodness-of-fit score to navigate the space of AGs, selecting an optimum as a representative for the MEC.
-Nonetheless, methods within both paradigms suffer from unreliability when data is scarce.
-Specifically, for the majority of the CD algorithms, formal assurances that the inferred MEC accurately represents the true causal model heavily rely on the so-called faithfulness assumption, which posits that all conditional independencies satisfied by $P(\mathbf{V})$ are entailed by the true causal model (Zhang and Spirtes, 2016).
-However, this presents a critical challenge in real-world scenarios, as violations of the faithfulness assumption become more prominent when relying on $P(\mathbf{V})$ estimated from limited data Uhler et al.
-(2012); Andersen (2013); Marx et al.
-(2021).
-For constraint-based methods, hypothesis tests may lack the statistical power to detect conditional independencies accurately.
-These errors may propagate and trigger a chain reaction of erroneous orientations Zhang and Spirtes (2008); Zhalama et al.
-(2017b); Ng et al.
-(2021).
-For score-based methods, although score functions directly measure goodness-of-fit on observational data, small sample sizes can significantly skew the estimates for the population parameters.
-Consequently, structures deemed score-optimal may not necessarily align with the ground-truth MEC  Ogarrio et al.
-(2016).
-A major concern is that the overwhelming majority of CD algorithms produce a single representation of the MEC as output, without quantifying the uncertainty that arises during the learning process Claassen and Heskes (2012); Jabbari et al.
-(2017).
-This poses a significant challenge for experts, as it hinders their ability to validate the algorithm’s outcome or gain insights into potential venues for improving inference quality.
-To alleviate the lack of uncertainty quantification in CD, we propose sampling AGs from a distribution defined using a score function, which places best-scoring AGs around the mode by design.
-This effectively provides end-users with samples that reflect the epistemic uncertainty inherent in CD, thus allowing their propagation through downstream causal inferences.
-In particular, we sample from our belief using Generative Flow Networks (GFlowNets; Bengio et al., 2021a, b), which are generative models known for sampling diverse modes while avoiding the mixing time problem of MCMC methods, and not requiring handcrafted proposals nor accept-reject steps (Bengio et al., 2021a).
-Acknowledging the low-data regime as CD’s primary challenge, we also propose actively integrating human feedback in the inferential process.
-This involves modeling user knowledge on the existence and nature (confounding/ancestral) of the relations and using it to weigh our beliefs over AGs.
-During our interactions with experts, we probe them about the relation of the pair of variables that optimizes a utility/acquisition function, for which we propose the negative expected cross-entropy between our prior and updated beliefs.
-Unlike prior strategies, our acquisition avoids the need to estimate the normalizing constant and predictive distribution of our updated belief, as needed for information gain and mutual information, respectively.
-Notably, we use importance sampling (Marshall, 1954; Geweke, 1989) to update our initial belief with the human feedback, which avoids retraining GFlowNets after each human interaction.
-While incorporating expert knowledge into CD has been a long-standing goal (Meek, 1995a; Chen et al., 2016; Li and Beek, 2018; Wang et al., 2022), existing works either rely on strong assumptions (e.g., causal sufficiency) or assume the knowledge is noiseless, aligned with the ground truth  (Andrews, 2020; Wang et al., 2022).
-Importantly, our work introduces the first iterative CD framework for AGs involving a human in the loop and accommodating potentially noisy feedback, as depicted in Figure 1.
-To validate our approach, we conduct experiments using the BIC score, for linear Gaussian causal models.
-Specifically, we assess: i) our ability to sample from score-based beliefs over AGs, ii) how our samples compare to samples from bootstrapped versions of state-of-the-art (SOTA) methods, and iii) the efficacy of our active knowledge elicitation framework using simulated humans.
-We observe that our method, Ancestral GFlowNet (AGFN), i) accurately samples from our beliefs over AGs; ii) consistently includes AGs with low structural error among its top-scored samples; and iii) is able to greatly improve performance metrics (i.e., SHD and BIC) when incorporating human in the loop.
-In summary, the contributions of our work are:
+Crucially, AGFN introduces a paradigm shift in causal inference by allowing the integration of uncertain knowledge in addition to the observational data.
+Incorporating background knowledge typically requires characterizing equivalence classes and designing specialized inference methods, but AGFN circumvents this by operating on distributions over AGs.
+Inference on each AG sample can be averaged, naturally propagating uncertainty.
+To validate our approach, we conduct experiments using the BIC score for linear Gaussian causal models, assessing i) our ability to accurately sample from score-based beliefs over AGs, ii) how these samples compare to those obtained from bootstrapped versions of state-of-the-art (SOTA) CD methods, and iii) the effectiveness of our active knowledge elicitation framework with simulated humans and an LLM.
+We observe that our method, AGFN, i) accurately samples from our beliefs over AGs; ii) consistently includes AGs with low structural error among its top-scored samples; and iii) is able to greatly improve performance metrics (i.e., SHD and BIC) when incorporating experts in the loop.
+In summary, our contributions are:
   1. 1.
-    We leverage GFlowNets to introduce AGFN, the first CD algorithm for scenarios under latent confounding that employs fully-probabilistic inference on AGs;
+    We propose AGFN, the first expert-in-the-loop probabilistic CD algorithm that supports uncertainty over the expert’s responses.
   2. 2.
-    We show AGFN accurately learns distributions over AGs, effectively capturing epistemic uncertainty.
+    We develop an elicitation framework that enables AGFN to optimally interact with experts by iteratively selecting maximally informative questions.
+This allows inference refinement that combines observational information with expert knowledge.
   3. 3.
-    We propose an experimental design to query potentially noisy expert insights on relationships among pairs of variables that lead to optimal uncertainty reduction.
+    We introduce a Bayesian algorithm to update AGFN’s distribution from expert responses, entirely bypassing model retraining.
   4. 4.
-    We show how to incorporate expert feedback into AGFN without retraining them from scratch.
+    We evaluate AGFN on diverse CD tasks, including the realistic benchmark Sachs dataset (Sachs et al.,, 2005).
+Results show that AGFN accurately and efficiently samples from its belief distribution and is competitive with SOTA CD in SHD and BIC.
+We also show that LLMs can be effectively used as surrogate experts for the EITL pipeline.
 2 BACKGROUND
 This section introduces the relevant notation and concepts.
-We use uppercase letters $V$ to represent a random variable or node in a graph, and boldface uppercase letters $\mathbf{V}$ to represent matrices or sets of random variables or nodes.
-Ancestral graphs.
-Under the assumption of no selection bias, an ancestral graph (AG) $\mathcal{G}$ over $\mathbf{V}$ is a directed graph comprising directed ($\rightarrow$) and bidirected ($\leftrightarrow$) edges Richardson and Spirtes (2002); Zhang (2007).
-In any directed graph, if a sequence of directed edges, $V_{i}\rightarrow\cdots\rightarrow V_{j}$, connects two nodes $V_{i}$ and $V_{j}$, we refer to this sequence as a directed path.
-In this case, we also say that $V_{i}$ is an ancestor of $V_{j}$ and denote this relation as $V_{i}\in An(V_{j})$.
+Uppercase letters ($V$) represent random variables or graph nodes.
+Bold uppercase letters ($\mathbf{V}$) represent matrices or sets of random variables or nodes.
+2.1 Ancestral graphs
+Assuming no selection bias, an ancestral graph (AG) $\mathcal{G}$ over $\mathbf{V}$ is a directed graph comprising directed ($\rightarrow$) and bidirected ($\leftrightarrow$) edges (Richardson and Spirtes,, 2002; Zhang,, 2007).
+In a directed graph, a sequence of directed edges $V_{i}\rightarrow\cdots\rightarrow V_{j}$ is called a directed path from $V_{i}$ to $V_{j}$.
+In this case, $V_{i}$ is an ancestor of $V_{j}$ and denote this relation as $V_{i}\in An(V_{j})$.
 By definition, any AG $\mathcal{G}$ must further satisfy the following:
@@
   2. 2.
-    there is no almost directed cycle, i.e., if $V_{i}\leftrightarrow V_{j}$ is in $\mathcal{G}$, then $V_{j}\not\in An(V_{i})$ and $V_{i}\not\in An(V_{j})$.
-As a probabilistic model, the nodes in an AG represent random variables, directed edges represent ancestral (causal) relationships, and bidirected edges represent associations solely due to latent confounding.
-For a complete characterization of AGs, refer to Richardson and Spirtes (2002).
-Data generating model.
-We assume that the data-generating model corresponds to a linear Gaussian structural causal model (SCM) (Pearl, 2000) defined by a 4-tuple $\mathcal{M}=\langle\mathbf{V},\mathbf{U},\mathcal{F},P(\mathbf{U})\rangle$, in which $\mathbf{V}=\{V_{1},V_{2},\dots,V_{n}\}$ is a set of $n$ observed random variables and $\mathbf{U}=\{U_{1},U_{2},\dots,U_{n}\}$ is the set of unobserved random variables.
+    there are no almost directed cycles, i.e., if $V_{i}\leftrightarrow V_{j}$ is in $\mathcal{G}$, then $V_{j}\not\in An(V_{i})$ and $V_{i}\not\in An(V_{j})$.
+As a probabilistic model, an AG has nodes as random variables, directed edges as ancestral (causal) relations, and bidirected edges as associations solely due to latent confounding (Richardson and Spirtes,, 2002).
+2.2 Linear Gaussian SCMs
+A linear Gaussian SCM is defined by a 4-tuple $\mathcal{M}=\langle\mathbf{V},\mathbf{U},\mathcal{F},P(\mathbf{U})\rangle$, in which $\mathbf{V}=\{V_{1},\dots,V_{n}\}$ is a set of $n$ observed random variables and $\mathbf{U}=\{U_{1},\dots,U_{n}\}$ is the set of unobserved random variables.
 Further, let $Pa_{i}\subseteq\mathbf{V}\setminus\{V_{i}\}$ be the set of observed causes (parents) of $V_{i}$, and $U_{i}$ be the set of unobserved causes of $V_{i}$.
-Then, each structural equation $f_{i}\in\mathcal{F}$ is defined as:
+Each structural equation $f_{i}\in\mathcal{F}$ is defined as:
 $V_{i}=\sum_{j:V_{j}\in Pa_{i}}\beta_{ij}V_{j}+U_{i}$ (1)
-with $P(\mathbf{U})$ being a multivariate Gaussian distribution with zero mean and a not necessarily identity covariance matrix $\boldsymbol{\Omega}=(\omega_{ij})_{1\leq i,j\leq n}$ — the error terms $\{U_{i}\}$ are not necessarily mutually independent, implying that the system can be semi-Markovian (i.e., latent confounding may be present).
-Consider a lower triangular matrix of structure coefficients $\mathbf{B}=(\beta_{ij})_{1\leq i,j\leq n}$ such that $(\mathbf{I}-\mathbf{B})$ is invertible, and $\beta_{ij}\neq 0$ only if $V_{j}\in Pa_{i}$.
-Then, the set of structural equations is given in matrix form by
+with $P(\mathbf{U})$ a multivariate Gaussian distribution with zero mean and a covariance matrix $\boldsymbol{\Omega}=(\omega_{ij})_{1\leq i,j\leq n}$ that is not necessarily the identity.
+The error terms $\{U_{i}\}$ are not necessarily mutually independent, implying that the system can include latent confounding.
+Let $\mathbf{B}=(\beta_{ij})_{1\leq i,j\leq n}$ be a lower-triangular matrix of structural coefficients such that $(\mathbf{I}-\mathbf{B})$ is invertible, with $\beta{ij}\neq 0$ only if $V_{j}\in Pa_{i}$.
+The structural equations can then be expressed in matrix form as
 $\mathbf{V}=\mathbf{B}\mathbf{V}+\mathbf{U}\implies\mathbf{V}=(\mathbf{I}-\mathbf{B})^{-1}\mathbf{U}.$ (2)
@@
 $\mathcal{N}_{\mathcal{M}}=\{\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma})|\boldsymbol{\Sigma}=(\mathbf{I}-\mathbf{B})^{-1}\boldsymbol{\Omega}(\mathbf{I}-\mathbf{B})^{-\top}\}$ (3)
-is represented by an AG in which, for every $i\neq j$, there is a directed edge $V_{j}\rightarrow V_{i}$ if $\beta_{ij}\neq 0$ and a bidirected edge $V_{j}\leftrightarrow V_{i}$ if $\omega_{ij}\neq 0$ (Richardson and Spirtes, 2002).
+is represented by an AG with directed edges $V_{j}\rightarrow V_{i}$ if $\beta_{ij}\neq 0$ and bidirected edges $V_{j}\leftrightarrow V_{i}$ if $\omega_{ij}\neq 0$, for every $i\neq j$ (Richardson and Spirtes,, 2002).
+To validate AGFN in our synthetic experiments, we draw samples from a linear Gaussian SCM and compare the AGFN-learned AG distribution against the ground truth.
 […]
-Figure 2: Illustration of the generative process of AGs $\{\mathcal{G}_{5},\mathcal{G}_{6},\mathcal{G}_{7}\}$ using GFlowNets.
-Starting with an empty graph $\mathcal{G}_{0}$, we add edges between variables $\{X_{1},X_{2},X_{3}\}$ according to the action-policy $\pi_{F}$.
-Solid edges trace trajectories leading to sampled graphs.
-Dashed lines represent non-realized transitions to terminal state $\square$.
-GFlowNets.
-Generative Flow Networks (GFlowNet; Bengio et al., 2021a, b) are generative models designed to sample from a finite domain $\mathcal{X}$ proportionally to some reward function $R:\mathcal{X}\rightarrow\mathbb{R}_{+}$, which may be parametrized using neural networks.
-In this work, we define $R$ as a strictly decreasing transformation of the BIC (more details in section 3).
-GFlowNets also assume there is a compositional nature to the elements $x\in\mathcal{X}$, meaning that they can be built by iteratively acting to modify a base object (i.e., an initial state).
-For instance, graphs can be built by adding edges to a node skeleton (Deleu et al., 2022) or molecules by adding atoms to an initial structure (Bengio et al., 2021a).
+Figure 2: Generative process of AGs $\{\mathcal{G}_{5},\mathcal{G}_{6},\mathcal{G}_{7}\}$ using GFlowNets.
+Starting from an empty graph $\mathcal{G}_{0}$, edges between variables $\{X_{1},X_{2},X_{3}\}$ are added following the action-policy $\pi_{F}$.
+Solid edges indicate trajectories leading to sampled graphs.
+Dashed lines denote non-realized transitions to the terminal state $\square$.
+2.3 GFlowNets
+GFlowNets are generative models designed to sample from a finite domain $\mathcal{X}$ proportionally to some reward function $R:\mathcal{X}\rightarrow\mathbb{R}_{+}$, which may be parametrized using neural networks.
+In this work, we define $R$ as a strictly decreasing transformation of the BIC (see more details in Section 3).
+GFlowNets also assume elements $x\in\mathcal{X}$ are compositional, built by iteratively modifying a base object (initial state).
+For instance, graphs can be generated by adding edges to an empty graph (Deleu et al.,, 2022), or molecules by adding atoms to an initial structure (Pandey et al.,, 2025).
 The generative process follows a trajectory of states $s\in\mathcal{S}$ guided by a transition probability $\pi_{F}:\mathcal{S}^{2}\rightarrow[0,1]$.
-In turn, $\pi_{F}$ is proportional to a foward flow function $F_{\theta}:\mathcal{S}^{2}\rightarrow\mathbb{R}_{+}$, which is parameterized by a neural network $\theta$.
+In turn, $\pi_{F}$ is proportional to a forward flow function $F_{\theta}:\mathcal{S}^{2}\rightarrow\mathbb{R}_{+}$, which is parametrized by a neural network $\theta$.
 Let $\text{Pa}(s^{\prime})$ ($\text{Ch}(s^{\prime})$) be the set of all states which can transition into (directly reached from) $s^{\prime}$.
@@
 There are also two special states in $\mathcal{S}$: an initial state $s_{0}$ and a terminal state $s_{f}$.
-We start with the initial state $s_{0}$ and transform it to a new valid state $s$ with probability $\pi_{F}(s|s_{0};\theta)$.
-We keep iterating this procedure until reaching $s_{f}$.
-States $s$ valid as final samples ($s\in\mathcal{X}$) are known as terminating states and have a positive probability for the transition $s\rightarrow s_{f}$.
+Starting from $s_{0}$, we iteratively transform it into a new valid state $s$ with probability $\pi_{F}(s\mid s_{0};\theta)$, repeating until reaching $s_{f}$.
+States $s$ that are valid as final samples ($s\in\mathcal{X}$) are known as terminating states and have a positive probability for the transition $s\rightarrow s_{f}$.
 Figure 2 illustrates this process with $\mathcal{X}$ being the space of AGs.
-Crucially, the same parameterization $\theta$ is used for all transition probabilities $\pi_{F}(\cdot|s;\theta)$ given any departing state $s$, allowing for generalization to states never visited during training.
-As the GFlowNet framework requires that no sequence of actions leads to a loop, we represent the space of possible action sequences by a pointed Directed Acyclic Graph (DAG) (Bengio et al., 2021b).
-The generation of any sample $x\in\mathcal{X}$ follows a trajectory $\tau=(s_{0},s_{1},\ldots,s_{T}=x,s_{f})\in\mathcal{S}^{T+2}$ for a $T\geq 0$.
-Different trajectories may lead to the same sample $x$.
-To ensure we sample proportionally to $R$, we search for a GFlowNet that satisfies the flow-matching condition, i.e., $\forall s^{\prime}\in\mathcal{S}$:
+The same parametrization $\theta$ is used for all transition probabilities $\pi_{F}(\cdot|s;\theta)$ given any departing state $s$, allowing for generalization to states not visited during training.
+As GFlowNets do not allow action sequences to form loops, the space of action sequences is represented as a pointed Directed Acyclic Graph (DAG) (Bengio et al., 2021b, ).
+The generation of any sample $x\in\mathcal{X}$ follows a trajectory $\tau=(s_{0},\ldots,s_{T}=x,s_{f})\in\mathcal{S}^{T+2}$ with $T\geq 0$.
+Many trajectories may reach a same $x$.
+To sample proportionally to $R$, we require a GFlowNet satisfying the flow-matching condition, i.e., $\forall s^{\prime}\in\mathcal{S}$:
 $\sum_{s\in\text{Pa}(s^{\prime})}F_{\theta}(s\rightarrow s^{\prime})=R(s^{\prime})+\!\!\sum_{s^{\prime\prime}\in\text{Ch}(s^{\prime})}\!\!F_{\theta}(s^{\prime}\rightarrow s^{\prime\prime}).$ (5)
-Equation 5 implies the flow that enters $s^{\prime}$ equals the flow leaving $s^{\prime}$, except for some flow $R(s^{\prime})$ leaking from $s^{\prime}$ into $s_{f}$.
-We let $R(s)=0$ for $s\notin\mathcal{X}$.
-Eventually, it may be that all states $s$ are valid candidates, i.e., $\mathcal{S}=\mathcal{X}\cup\{s_{f}\}$.
-If so, each of eq.
-5’s solutions satisfies a detailed-balance condition,
+Eq.
+(5) states that the flow into $s^{\prime}$ equals the flow out of $s^{\prime}$, except for some flow $R(s^{\prime})$ leaking from $s^{\prime}$ into $s_{f}$, with $R(s)=0$ for $s\notin\mathcal{X}$.
+When all states are valid samples, i.e., $\mathcal{S}=\mathcal{X}\cup\{s_{f}\}$, every solution of Eq.(5) satisfies a detailed-balance condition,
 $\frac{R(s)F_{\theta}(s\rightarrow s^{\prime})F_{\theta}(s^{\prime}\rightarrow s_{f})}{F_{\theta}(s\rightarrow s_{f})}=R(s^{\prime})F_{B,\theta}(s^{\prime}\rightarrow s),\!$ (6)
-for a parametrized backward flow $F_{B,\theta}\colon\mathcal{S}^{2}\rightarrow\mathbb{R}_{+}$ Deleu et al. (2022). In practice, we enforce eq. 6 by minimizing
+for some backward flow $F_{B,\theta}\colon\mathcal{S}^{2}\rightarrow\mathbb{R}_{+}$ (Deleu et al.,, 2022). In practice, we enforce (6) by minimizing
 $\mathcal{L}(\theta)\!\!=\!\!\underset{s\rightarrow s^{\prime}}{\mathbb{E}}\!\left[\!\left(\!\log\frac{R(s^{\prime})\pi_{F_{B}}(s|s^{\prime};\theta)\pi_{F}(s_{f}|s;\theta)}{R(s)\pi_{F}(s^{\prime}|s;\theta)\pi_{F}(s_{f}|s^{\prime};\theta)}\!\right)^{2}\!\right].$ (7)
 3 ANCESTRAL GFLOWNETS
-We propose AGFN, a GFlowNet-based method for sampling AGs using a score function. Specifically, AGFN encompasses a GFlowNet with the following characteristics:
+To quantify uncertainty in CD under latent confounding and enhance inferences via active learning, we propose AGFN, a GFlowNet-based probabilistic method that samples AGs from a score-based belief distribution.
+AGFN is built around a GFlowNet in which:
   1. 1.
@@
     A well-trained AGFN samples AGs with frequencies proportional to their rewards and with the best-scoring AG being, by design, the mode.
-[…]
-Action constraints.
-To ensure AGFN only samples AGs, we mask out actions that would lead to paths forming cycles or almost cycles.
-To achieve this, we verify whether the resulting graph respects Bhattacharya et al.
-(2021)’s algebraic characterization of the space of AGs.
-More specifically, any AG $\mathcal{G}$ is characterized by an adjacency matrix $\mathbf{A}_{d}\in\mathbb{R}^{n\times n}$ for its directed edges and another adjacency matrix $\mathbf{A}_{b}\in\mathbb{R}^{n\times n}$ for its bidirected edges, adhering to:
-$\text{trace}(\exp\{\mathbf{A}_{d}\})-n+\mathbf{1}^{T}(\exp\{\mathbf{A}_{d}\}\odot\mathbf{A}_{b})\mathbf{1}=0,$ (8)
-in which $\mathbf{1}$ is a $n$-dimensional unit vector and $\odot$ denotes the Hadamard (elementwise) product of matrices.
-Score-based belief. We propose using a strictly decreasing transformation of a score function $U$ as the reward $R$ for AGFN. More precisely, we define $R$ as
-$R(\mathcal{G})=\exp\left\{\frac{\mu-U(\mathcal{G})}{\sigma}\right\}$ (9)
-for given constants $\mu\in\mathbb{R}$ and $\sigma\in\mathbb{R}^{+}$ that ensure numerical stability (Zhang et al., 2023).
-In practice, we sample $S$ AGs $\{\mathcal{G}^{(s)}\}_{s=1}^{S}$ from an untrained AGFN, and set $\mu=\nicefrac{{1}}{{S}}\sum_{s}U(\mathcal{G}^{(s)})$ and $\sigma=\sqrt{\nicefrac{{1}}{{S}}\sum_{s}(U(\mathcal{G}^{(s)})-\mu)^{2}}$.
-Score for linear Gaussian models.
-Since we focus on linear Gaussian models, we choose the extended Bayesian Information Criterion Foygel and Drton (2010) as our score function.
-Specifically, for any AG $\mathcal{G}=(\mathbf{V},\mathbf{E})$:
-$U(\mathcal{G})$ $=-2l_{N}(\hat{\mathbf{B}},\hat{\boldsymbol{\Omega}})+|\mathbf{E}|\log n+2|\mathbf{E}|\log|\mathbf{V}|,$ (10)
-in which $(\hat{\mathbf{B}},\hat{\boldsymbol{\Omega}})$ is the MLE estimate of model parameters (see eq. 3) obtained using the residual iterative conditional fitting algorithm Drton et al. (2009).
-Forward flow.
-We use a Graph Isomorphism Network (GIN) Xu et al.
-(2019) $\Phi$ to compute a $d$-dimensional representation for each node in the AG $\mathcal{G}_{t}$ at the $t$-th step of the generative process and use sum pooling to get an embedding for $\mathcal{G}_{t}$.
-Then, considering $\mathcal{A}_{t}$ as the space of feasible actions at $\mathcal{G}_{t}$ (i.e., those leading to an AG), we use an MLP $\phi\colon\mathbb{R}^{d}\rightarrow\mathbb{R}^{|\mathcal{A}_{t}|}$ with a softmax activation at its last layer to map $\mathcal{G}_{t}$’s embedding to a distribution over $\mathcal{A}_{t}$.
-More precisely, given $\mathbf{H}^{(t)}=\Phi(\mathcal{G}_{t})\in\mathbb{R}^{|\mathbf{V}|\times d}$, we compute
-$\mathbf{p}=\phi\left(\sum_{v\in\mathbf{V}}\mathbf{H}_{v}^{(t)}\right)$ (11)
-as the probability distribution over the feasible actions at $\mathcal{G}_{t}$.
-Backward flow.
-Backward actions correspond to removing edges.
-Following Shen et al.
-(2023), we parametrize the backward flow $F_{B}$ with an MLP and alternate between updating $\pi_{F}$ and $\pi_{F_{B}}$, using gradient-based optimization.
-4 Human-in-the-Loop Causal Discovery
-Concluding the training, we propose leveraging the AGFN-generated samples to design questions for the expert that optimize the reduction of entropy in the distribution $p_{\theta}(\mathcal{G})$ over the space of AGs.
-Then, we use the human feedback to update $p_{\theta}(\mathcal{G})$, and iteratively repeat the process.
-The following paragraphs describe i) how we model human feedback, ii) how we update our belief over AGs given human responses, and iii) our experimental design strategy for expert inquiry.
-Modeling human feedback.
-We assume humans are capable of answering questions regarding the ancestral relationship between pairs of random variables.
-In this case, we model their prior knowledge on a relation $r=\{U,V\}$ between nodes $U,V\in\mathbf{V}$ as a categorical distribution over a random variable denoted $\omega_{r}$.
-Fix an arbitrary total order $<$ in $\mathbf{V}$.
-By definition, $\omega_{r}=1$ if there is no edge between $U$ and $V$; $\omega_{r}=2$ if $U$ is ancestor of $V<U$; $\omega_{r}=3$ if $V$ is ancestor of $U>V$; and $\omega_{r}=4$ if there is a bidirected edge between $U$ and $V$.
-Since the human has access to our AGFN before being probed for the first time, we set $\rho_{r,k}=p_{\theta}(\omega_{r}=k)$ as the prior probability of $\omega_{r}=k$.
-Moreover, we consider that the expert’s feedback $f_{r}\in\{1,2,3,4\}$ on the relation $r$ is a noisy realization of the true, unobserved value of the relation feature $\omega_{r}$ under the expert’s model.
-Putting all elements together results in a two-level Bayesian hierarchical scheme for categorical data:
+The generation of a trajectory $\tau=\{\{\},\mathcal{G}_{1},\mathcal{G}_{2},\ldots,\mathcal{G}_{T}\}$ begins with an empty graph with nodes $\mathbf{V}$, iteratively adding edges of types $\{\leftarrow,\rightarrow,\leftrightarrow\}$ between pairs of variables.
+The following sections describe AGFN.
+For additional details refer to the Appendix.
+3.1 Action constraints
+To ensure AGFN samples only valid AGs, we restrict actions that would create cycles or almost cycles.
+An action adding a directed edge leads to a cycle (resp.
+almost cycle) if and only if a directed path (resp.
+directed path with a single bidirected edge) exists between its endpoints, with a similar reasoning for adding bidirected edges.
+Similarly to Giudici and Castelo, (2003) and (Deleu et al.,, 2022, Appendix C), we leverage the iterative nature of the generative process to efficiently track such paths and mask invalid actions, as discussed in Appendix A.
+3.2 Score-based belief
+We define AGFN’s reward $R$ as a strictly increasing transformation of a score function $U$ measuring the goodness-of-fit of an AG to the observed data.
+To ensure numerical stability (Zhang et al.,, 2023), we choose constants $\mu\in\mathbb{R}$ and $\sigma\in\mathbb{R}^{+}$ and define the reward function $R(\cal{G})$ for the AG $\cal{G}$ as
+$R(\mathcal{G})=\exp\left\{\frac{U(\mathcal{G})-\mu}{\sigma}\right\}.$ (8)
+In practice, we sample $S$ AGs, $\{\mathcal{G}^{(s)}\}_{s=1}^{S}$, from an untrained AGFN, and define
+$\mu=\frac{1}{S}\sum_{s}U(\mathcal{G}^{(s)});\,\,\sigma=\sqrt{\frac{1}{S}\sum_{s}(U(\mathcal{G}^{(s)})-\mu)^{2}}.$ (9)
+Drawing on the generative process described in Section 3.1 and the reward function defined above, we learn an AGFN by minimizing the loss function in Equation (7) through stochastic gradient descent (Kingma and Ba,, 2017).
+For further details regarding AGFN training and the parametrization of the GFlowNet’s policies, please consult Section C.2.1.
+4 EITL Causal Discovery
+After training, we leverage AGFN-generated samples to design expert queries and refine the learned distribution based on their answers.
+Queries are selected to maximally reduce the entropy of the distribution $p_{\theta}(\mathcal{G})$ over AGs.
+Expert feedback is then used to update $p_{\theta}(\mathcal{G})$, and the process is repeated.
+Next we describe i) how feedback is modeled, ii) how beliefs over AGs are updated given expert responses, and iii) our experimental design strategy for expert queries.
+4.1 Modeling expert feedback
+We leverage expert feedback to determine the existence of ancestral relationships between nodes.
+A detailed discussion of this type of knowledge is in Section 4.5.
+We model prior knowledge on a relation $r=\{V_{i},V_{j}\}$ between nodes $V_{i},V_{j}\in\mathbf{V}$ as a categorical distribution over a random variable denoted $\omega_{r}$.
+For notational convenience, we let $\omega_{r}=1$ if there is no edge between $V_{i}$ and $V_{j}$; $\omega_{r}=2$ if $V_{i}$ is ancestor of $V_{j}<V_{i}$; $\omega_{r}=3$ if $V_{j}$ is ancestor of $V_{i}>V_{j}$; and $\omega_{r}=4$ if there is a bidirected edge between $V_{i}$ and $V_{j}$.
+Any other encoding for $\omega_{r}$ would yield an equivalent algorithm.
+Since the expert sees AGFN before being the first query, we set $\rho_{r,k}=p_{\theta}(\omega_{r}=k)$ as the prior probability of $\omega_{r}=k$.
+The expert’s feedback $f_{r}\in\{1,2,3,4\}$ is treated as a noisy realization of the true, unobserved relation feature $\omega_{r}$.
+Together, these elements form a Bayesian hierarchical scheme for categorical data:
 $\omega_{r}$ $\sim\text{Cat}(\boldsymbol{\rho}_{r}),$ (10)
@@
 $\frac{\boldsymbol{\rho}_{r}}{\eta_{r}}\odot\left(\pi\cdot\delta_{f_{r}}+\left(\frac{1-\pi}{3}\right)\cdot(\mathbf{1}-\delta_{f_{r}})\right),$ (12)
-with $\eta_{r}=\rho_{r,f_{r}}\cdot\pi+\left(\frac{1-\pi}{3}\right)\cdot(1-\rho_{r,f_{r}})$.
-Figure 3: Sampling quality. The reward-induced marginal distribution over graph features is adequately approximated by the marginal distribution learned by the GFlowNet.
-Updating beliefs.
-We update our AGFN by weighing it by our posterior over the expert’s knowledge, described in the previous paragraph, similarly to a product-of-experts approach Hinton (2002).
-For this, let $\mathbf{f}_{K}=(f_{r_{k}})_{1\leq k\leq K}$ be the sequence of $K$ feedbacks issued by the expert and define our novel belief distribution $q(\mathcal{G};\mathbf{f}_{K})$ over the space of AGs as
-$q(\mathcal{G};\mathbf{f}_{K})\propto p_{\theta}(\mathcal{G})\prod_{1\leq k\leq K}p(\omega_{r_{k}}|f_{r_{k}}).$ (15)
-Importantly, we use $p_{\theta}$ as proposal distribution in importance (re-)sampling Gordon et al.
-(1993) to approximately sample from $q(\mathcal{G};\mathbf{f}_{K})$ — or to approximate the expected value of a test function.
-More precisely, we estimate the value of a function $h$ over the space of AGs as:
-$\mathbb{E}_{q}[h(\mathcal{G})]\approx\sum_{t=1}^{T}c^{-1}\frac{q(\mathcal{G}^{(t)};\mathbf{f}_{K})}{p_{\theta}(\mathcal{G}^{(t)})}h(\mathcal{G}^{(t)})$ (16)
-with $(\mathcal{G}^{(t)})_{t=1}^{T}\sim p_{\theta}(\mathcal{G})$ and $c=\sum_{t=1}^{T}\nicefrac{{q(\mathcal{G}^{(t)};\mathbf{f}_{K})}}{{p_{\theta}(\mathcal{G}^{(t)})}}$.
-Active knowledge elicitation.
-To make the most out of possibly costly human interactions, we query the human about the relation that maximally reduces the expected cross-entropy between our belief over AGs before and after human feedback.
-More precisely, we define an acquisition function $a_{k}:{{\mathbf{V}}\choose{2}}\rightarrow\mathbb{R}$ for the $k>1$-th inquiry as:
-$a_{k}(r)=-\mathbb{E}_{f_{r}\sim p(\cdot|\mathbf{f}_{K})}\big{[}\mathbf{H}\left(q(\mathcal{G};\mathbf{f}_{K},f_{r}),q(\mathcal{G};\mathbf{f}_{K})\right)\!\big{]}$ (17)
-in which $p(f_{r}|\mathbf{f}_{K})$ is the posterior predictive distribution according to the user model, $q_{0}\propto R$ and $\mathbf{H}(\cdot,\cdot)$ is the cross-entropy.
-Then, we maximize this acquisition to select which relation $\tilde{r}_{k}$ we will probe the expert, i.e.:
-$\tilde{r}_{k}=\operatorname*{arg\ max}_{r\in{{\mathbf{V}}\choose{2}}}a_{k}(r).$ (18)
-As aforementioned, we use importance sampling with $q_{0}$ as a proposal to estimate the acquisition function $a_{k}$.
-This allows us to leverage AGFN samples and effectively avoid the need for retraining them.
-It is worth mentioning that because $\mathbf{H}(p,p^{\prime})\geq\mathbf{H}(p,p)$ for any two distributions $p$ and $p^{\prime}$ of the same support, our strategy is equivalent to minimizing an upper bound on the entropy of $q_{k}$.
-Also, different from acquisitions based on information gain or mutual information, approximating Equation 17 via Monte Carlo does not require exhaustive integration over the space of AGs to yield asymptotically unbiased estimates — see Appendix.
-chain4 IV collfork
-SHD BIC SHD BIC SHD BIC
-FCI $3.03{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.13}}$ 1.13 $5481.33{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.69}}$ 2.69 $\mathbf{3.75}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 0.64}}$ 0.64 $\mathbf{5426.18}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.74}}$ 1.74 $6.26{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.20}}$ 1.20 $5433.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 6.94}}$ 6.94
-GFCI $\mathbf{2.24}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 0.64}}$ 0.64 $\mathbf{5479.77}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.75}}$ 1.75 $4.21{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 0.96}}$ 0.96 $5427.09{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.85}}$ 2.85 $\mathbf{5.23}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.08}}$ 1.08 $\mathbf{5431.67}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 7.91}}$ 7.91
-DCD $3.38{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.30}}$ 1.30 $5482.97{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 5.16}}$ 5.16 $5.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.23}}$ 1.23 $5429.51{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 4.37}}$ 4.37 $6.02{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.22}}$ 1.22 $5436.84{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 9.41}}$ 9.41
-[…]
-AGFN (ours) $\mathbf{6.04}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.12}}$ 2.12 $\mathbf{5494.67}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 37.08}}$ 37.08 $8.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.04}}$ 2.04 $\mathbf{5456.16}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 52.25}}$ 52.25 $\mathbf{6.58}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.34}}$ 2.34 $\mathbf{5478.01}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 40.36}}$ 40.36
-Table 1: Average SHD and BIC.
-The denotes methods yielding point estimates.
-We use Bootstrap to report the mean and average standard deviation for these.
-For N-ADMG and AGFN, we estimate the quantities using $100$k samples.
+with $\eta_{r}=\rho_{r,f_{r}}\cdot\pi+\left(\frac{1-\pi}{3}\right)\cdot(1-\rho_{r,f_{r}})$ and $\odot$ representing the Hadamard product between vectors.
+Figure 3: Sampling quality.
+AFGN accurately samples from its underlying score-based beliefs.
+The 1st plot (left) shows that the marginal probabilities over edges induced by AGFN match the analytical marginal.
+The 2nd shows the same for the probability of directed paths between two variables.
+The rightmost plots show that the distributions of the SHD and BIC from AGFN samples closely match the analytical one.
+4.2 Updating beliefs
+To understand our procedure for updating AGFN’s distribution based on the observed responses from an expert, the reader is invited to notice that the expert-induced posterior, $p(f_{r}|w_{r})$, characterizes a distribution over the edges of an AG through $q(e|\mathbf{f}_{K})\coloneqq p(w_{r}|f_{r})\cdot u(r)$.
+Here, $q(e)$ denotes the probability of an edge $e$ corresponding to the relation $r$ featuring $f_{r}$ (as described in Section 4.1) and $u(r)$ is an uniform distribution over the AGs’ unlabeled edges.
+Similarly, an AGFN’s policy $\pi_{F}(\cdot|s)$ corresponds to a distribution over the edges of the AG $s$, which we denote by $p(e)$.
+From this perspective, we follow a mixture-of-experts approach (Hinton,, 2002) and let the updated AGFN’s policy be $\pi_{F}^{(K)}(s^{\prime}|s,\mathbf{f}_{K})\propto p(e)\cdot q(e|\mathbf{f}_{K})$, in which $e$ is the edge added in the transition from $s$ to $s^{\prime}$.
+Intuitively, this ensures that the AGFN samples graphs whose edges are consistent with both the score function and the expert’s beliefs more frequently.
+Additionally, the structural constraints for ancestrality of the sampled graphs remain naturally encoded into $p(e)$ (see Figure 2).
+Based on the updated policy, we can easily estimate the expectation of a given test function (e.g., an acquisition function for active learning; see Section 4.3) through a Monte Carlo estimator.
+4.3 Active knowledge elicitation
+To efficiently utilize possibly costly expert interactions, we query the relation that maximally reduces the expected cross-entropy between our belief over AGs before and after feedback.
+Specifically, we define an acquisition function $a_{k}:{{\mathbf{V}}\choose{2}}\rightarrow\mathbb{R}$ for the $k>1$-th inquiry as:
+$a_{k}(r)=-\mathbb{E}_{f_{r}\sim p(\cdot|\mathbf{f}_{K})}\big[\mathbf{H}\left(q(\mathcal{G};\mathbf{f}_{K},f_{r}),q(\mathcal{G};\mathbf{f}_{K})\right)\!\big],$ (13)
+in which $p(f_{r}|\mathbf{f}_{K})$ is the posterior predictive distribution under the user model, $q_{0}\propto R$, and $\mathbf{H}(\cdot,\cdot)$ is the cross-entropy.
+We then maximize this acquisition to select the relation $\tilde{r}_{k}$ to query the expert, i.e.:
+$\tilde{r}_{k}=\operatorname*{arg\ max}\nolimits_{r\in{{\mathbf{V}}\choose{2}}}a_{k}(r).$ (14)
+Since $\mathbf{H}(p,p^{\prime})\geq\mathbf{H}(p,p)$ for any distributions $p$ and $p^{\prime}$ of the same support, our strategy is equivalent to minimizing an upper bound on the entropy of $q_{k}$.
+Unlike acquisitions based on information gain or mutual information, Monte Carlo approximation of Eq.
+(13) avoids exhaustive integration over the space of AGs to yield asymptotically unbiased estimates (see Appendix C).
+Algorithm 1 Simulating experts in the loop
+$\{\mathcal{G}_{t}\}_{1\leq t\leq T}\text{ samples from AGFN}$, $\mathcal{G}^{*}=(\mathbf{V},E)$ true AG, $\pi$ feedback reliability
+$p(\omega_{r}=k)\leftarrow\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}\{\omega_{r}=k\text{ in }\mathcal{G}_{t}\},\forall k\in\{1,2,3,4\},\ r\in{\mathbf{V}\choose 2}$
+$\mathbf{f}\leftarrow\{\}$ $\triangleright$ Set of feedbacks (answers)
+$\boldsymbol{r}\leftarrow\{\}$ $\triangleright$ Set of queries (questions)
+$K\leftarrow 1$
+$\omega_{r}^{*}\leftarrow\text{ relation }r\text{'s feature in }\mathcal{G}^{*}\ \forall r\in{\mathbf{V}\choose 2}$
+while $\boldsymbol{r}\neq{\mathbf{V}\choose 2}$ do $\triangleright$ Iteratively query expert
+  $r_{K}\leftarrow\underset{r\in{\mathbf{V}\choose 2}\setminus\boldsymbol{r}}{\operatorname*{arg\ max}}\ \underset{{f_{r}\sim p(\cdot)}}{\mathbb{E}}[-\mathbf{H}\left(q(\mathcal{G};\mathbf{f}\cup\{f_{r}\}),q(\mathcal{G};\mathbf{f})\right)]$
+  $\mathbf{r}\leftarrow\mathbf{r}\cup\{r_{K}\}$
+  $f_{K}\sim\text{Cat}\left(\pi\cdot\delta_{\omega_{r_{K}}^{*}}+\left(\frac{1-\pi}{3}\right)\cdot(1-\delta_{\omega_{r_{K}}^{*}})\right)$
+  $\mathbf{f}\leftarrow\mathbf{f}\cup\{f_{K}\}$
+  $K\leftarrow K+1$
+end while
+4.4 Algorithmic details
+Algorithm 1 outlines our procedure for simulating expert interactions.
+We first estimate the marginal probabilities $p(\omega_{r}=k)$ of a relation $r$ having feature $k\in\{1,2,3,4\}$ under AGFN’s learned distribution, which serves as our prior distribution.
+We then iteratively select the relation that maximizes the acquisition function; the expert returns the relation’s true feature with probability $\pi$, or otherwise chooses randomly among the incorrect options.
+4.5 On (non-)ancestral knowledge
+Our EITL approach leverages the fact that experts can often easily judge whether one variable is an ancestor of another, without needing to determine whether the causal link is direct or confounded.
+Such coarse knowledge is easier to provide than the direct causation required for DAGs.
+For example, experimental studies frequently establish ancestrality between variables, even if they cannot rule out mediation or confounding.
+Likewise, experts may know that demographic or genetic variables are not downstream of diseases, or that certain associations are spurious.
+Our framework integrates such knowledge, along with the expert’s confidence level, by appropriately adjusting the probabilities of ancestral, non-ancestral, or confounding relationships.
+A detailed discussion is in Appendix F.
 5 EXPERIMENTS
 Our experiments have three objectives.
-First, we validate that AGFN can accurately learn the target distribution over the space of AGs.
-Second, we show that AGFN performs competitively with alternative methods on three data sets.
+First, we validate that AGFN approximates well the target distribution over the space of AGs.
+Second, we show that AGFN performs competitively with alternative methods on both synthetic and real-world datasets.
 Third, we attest that our experimental design for incorporating the expert’s feedback efficiently reduces the uncertainty over AGFN’s distribution.
-We provide further experimental details in the Appendix.
-Code is available in the supplement.
+We provide further experimental details in Appendix C, jointly with results on different datasets, such as DREAM3 Challenge (Marbach et al.,, 2009).
+Code for reproducing the results were submitted as supplemental material.
 5.1 Distributional Assessment of AGFN
 Data.
-Since violations of faithfulness are more likely in dense graphs Uhler et al.
-(2012), we create $20$ $5$-node random graphs Uhler et al.
-(2012) from a directed configuration model (Newman, 2010) whose in- and out-degrees are uniformly sampled from $\{0,1,2,3,4\}$.
-We draw $500$ independent samples from a structure-compatible linear Gaussian SCM with random parameters for each graph.
+As faithfulness violations are more likely in dense graphs (Uhler et al.,, 2012), we generate 2020 55-node random graphs using a directed configuration model (Newman,, 2010), with in- and out-degrees uniformly sampled from $\{0,\ldots,4\}$.
+For each graph, we draw $500$ samples from a structure-compatible linear Gaussian SCM with randomly chosen parameters.
+Setup.
+We train AGFN for each random graph using their respective samples.
+Then, we collect AGFN samples to compute empirical distributions over the i) edge features (i.e., $p_{\theta}(V_{i}\rightarrow V_{j})$ and $p_{\theta}(V_{i}\leftrightarrow V_{j})$ for each pair $(V_{i},V_{j})$), ii) path features (i.e., $p_{\theta}(V_{i}\rightsquigarrow V_{j})$) iii) BIC, and iv) Structural Hamming Distance (SHD) to the true causal diagram.
+Results.
+Figure 3 shows that AGFN adequately approximates the theoretical distribution induced by Equation 8.
+It also induces BIC and SHD distributions that closely match those of $p(\mathcal{G})\propto R(\mathcal{G})$.
+Here we highlight a critical limitation of N-ADMG, a baseline probabilistic CD method: over $60\%$ of its generated graphs were non-ancestral, severely limiting its utility for inference over AGs.
+By contrast, AGFN ensures that all generated samples are valid AGs.
+5.2 Comparison with SOTA CD algorithms
+Data.
+We generate $10$ datasets with $500$ independent samples from the randomly parametrized linear Gaussian SCMs corresponding to the canonical causal diagram (Richardson and Spirtes,, 2002) of four 4-node AGs with increasingly structural complexity: i) chain4: $W\rightarrow X\rightarrow Y\rightarrow Z$, a simple chain without latent confounders; ii) collfork: $X\rightarrow Z\leftarrow W\rightarrow Y$; with $X\leftrightarrow W$ and $Z\leftrightarrow Y$, forming triplets with colliders and non-colliders under latent confounding, and iii) IV: $X\rightarrow Y$ with $W\rightarrow X\leftarrow Z\rightarrow Y$, a discriminating path for $Z$.
+Those capture unshielded colliders and discriminating paths, fundamental patterns identifiable by CD algorithms under latent confounding (Spirtes and Richardson,, 1997; Zhang, 2008b, ).
+Figure 4: Human-aided AGFN outperforms CD baselines after a single feedback in the considered datasets. Each plot shows the expected SHD under a varying number of feedbacks over $30$ EITL simulations.
+Baselines.
+We compare AGFN with five notable CD methods: FCI (Spirtes et al.,, 2001; Zhang, 2008b, ), GFCI (Ogarrio et al.,, 2016), ACI (Magliacane et al.,, 2016), DCD (Bhattacharya et al.,, 2021), and N-ADMG (Ashman et al.,, 2023), spanning different CD paradigms.
+FCI is a seminal constraint-based CD algorithm that learns a PAG consistent with conditional independencies entailed by statistical tests.
+GFCI is a hybrid CD algorithm that learns a PAG by first obtaining an approximate structure using FGS (Ramsey,, 2015) (a BIC-score-based search algorithm for causally sufficient scenarios) and then by applying FCI to identify possible confounding and remove some edges added by FGS.
+ACI (Magliacane et al.,, 2016) translates conditional independencies into weighted logical statements and uses an Answer Set Programming solver to find an optimal AG.
+DCD casts CD as continuous optimization with differentiable algebraic constraints defining the space of AGs and uses gradient-based algorithms to solve it.
+N-ADMG computes a variational approximation of the joint posterior distribution over the space of bow-free causal diagrams (Nowzohour et al.,, 2017) of non-linear SCMs with additive noise.
+For a more detailed description, we refer to Section C.1.
+Experimental setup.
+We train AGFN on each dataset and sample 100k AGs.
+For comparison, we apply FCI, GFCI, ACI, and DCD to 100 bootstrapped resamplings of each dataset to approximate the confidence distributions these algorithms induce.
+Also, we compute the sample mean and standard deviation of BIC and SHD between each method’s estimated PAG and the ground-truth PAG.
+For methods that output a PAG member (DCD, N-ADMG, and AGFN) we first extract the corresponding PAG before computing the SHD.
+BIC, is computed directly, as all PAG members are asymptotically score-equivalent.
+Results.
+Summary statistics for the distribution over AGs induced by AGFN and by other baseline CD algorithms are provided in Table 4 in the supplement.
+Notably, both probabilistic and non-probabilistic methods achieve comparable BIC performance, with AGFN incurring in significantly more diverse samples.
+Table 1 shows that the three highest-reward samples from AGFN match or exceed the performance of the baselines.
+For N-ADMG, results comprise the three most frequent samples from its variational distribution.
+Table 1: SHD for point estimates. Mean SHD of the AGFN’s top-3 draws is comparable to or better than baselines.
+chain4 IV collfork
 […]
+[…]
+[…]
+[…]
+[…]
+[…]
+5.3 Simulating experts in the loop
+Data.
+Following the procedure from Section 5.1, we generate graphs with $4$, 66, 88 and $10$ nodes and draw $500$ samples from a compatible linear Gaussian SCM to train AGFN.
+We then apply our active elicitation strategy (Alg.
+1) to probe simulated experts, adhering to the generative model in Section 4, with $\pi=0.9$.
+Setup.
+Since benchmarks for expert knowledge elicitation are lacking, we assess whether incorporating expert feedback enhances the concentration of the learned distribution around the true AG and evaluate our elicitation strategy by tracking SHD and BIC as a function of the number of expert interactions.
 Results.
-Figure 3 shows that the AGFN adequately approximates the theoretical distribution induced by the reward in Equation 9.
-Furthermore, AGFNs induce distributions over BIC and SHD values that closely resemble those induced by $p(\mathcal{G})\propto R(\mathcal{G})$.
-We also note an important improvement over the prior art on probabilistic CD (N-ADMG): we found that over $60\%$ of its samples were non-ancestral and that this method was of little use for making inferences over AGs.
-Meanwhile, AGFN does not sample non-ancestral graphs.
-5.2 Comparison with SOTA CD algorithms
-(a) chain4
-(b) IV
-(c) collfork
-Figure 4: Ancestral graphs representing the data generating models for the three considered datasets in Table 1.
+Figures 4 and 5 show that the expected SHD under our belief over AGs substantially decreases with expert feedback.
+The consistent decrease in SHD indicates that our belief increasingly concentrates on the true AG as feedback is iteratively incorporated, regardless of the querying strategy.
+Moreover, our querying strategy accelerates BIC reduction compared to random queries (see Section C.8), confirming that some edges are more informative and should be prioritized when querying the expert.
+Expert-aided AGFN versus baselines.
+Figure 4 indicates that updating AGFN’s distribution by incorporating the expert’s feedback leads to a significant enhancement of the model’s accuracy in finding the underlying causal diagram (ACI is omitted due to its significant lower performance).
+Remarkably, even a single feedback is sufficient for AGFN to outperform all considered baselines.
+Importantly, integrating expert knowledge into prior CD methods is generally nontrivial: when feasible, it typically occurs before inference (Andrews,, 2020) or assumes perfect rather than uncertain knowledge (Wang et al.,, 2022).
+5.4 GPT-4o as the expert-in-the-loop
+Figure 5: GPT-aided AGFN outperforms CD baselines in the considered datasets. Each plot shows the expected SHD under a varying number of feedbacks over 33 EITL simulations.
 Data.
-We generate $10$ datasets with $500$ independent samples from the randomly parametrized linear Gaussian SCMs corresponding to the canonical causal diagram Richardson and Spirtes (2002) in each AG depicted in Figure 4.
-Unshielded colliders and discriminating paths are fundamental patterns in the detection of invariances by CD algorithms under latent confounding Spirtes and Richardson (1997); Zhang (2008b).
-Thus, we consider the following 4-node causal diagrams with increasingly difficult configurations: (i) chain4, a chain without latent confounders; (ii) collfork, a graph with triplets involving colliders and non-colliders under latent confounding, and (iii) IV, a structure with a discriminating path for $Z$: $W\rightarrow X\leftarrow Z\rightarrow Y$.
-Baselines.
-We compare AGFN with four notable CD methods: FCI (Spirtes et al., 2001; Zhang, 2008b), GFCI (Ogarrio et al., 2016), DCD (Bhattacharya et al., 2021), and N-ADMG Ashman et al.
-(2023).
-The baselines span four broad classes of CD methods.
-FCI is a seminal constraint-based CD algorithm that learns a PAG consistent with conditional independencies entailed by statistical tests.
-GFCI is a hybrid CD algorithm that learns a PAG by first obtaining an approximate structure using FGS (Ramsey, 2015) (a BIC-score-based search algorithm for causally sufficient scenarios) and then by applying FCI to identify possible confounding and remove some edges added by FGS.
-DCD casts CD as continuous optimization with differentiable algebraic constraints defining the space of AGs and uses gradient-based algorithms to solve it.
-N-ADMG computes a variational approximation of the joint posterior distribution over the space of bow-free causal diagrams (Nowzohour et al., 2017) associated with non-linear SCMs with additive noise.
-While N-ADMG focuses on a more restricted setting compared to AGs, it offers some uncertainty quantification in the variational pohttps://www.overleaf.com/project/650454a084b798332af29ebesterior, making it more closely comparable to our approach.
-We rigorously follow the experimental guidelines in the original works.
-Experimental setup We train AGFN on each dataset and use it to sample 100k graphs.
-We also apply FCI, GFCI, and DCD to $100$ bootstrapped resamplings of each dataset to emulate confidence distributions induced by these algorithms.
-To compare the algorithms’ outputs, we compute the sample mean and standard deviation of the BIC and SHD at the PAG level.
-Specifically, we compute the SHD between the ground-truth PAG and each estimated PAG obtained by each method.
-If the output is a PAG member (as for DCD, N-ADMG, and AGFN) we use FCI to transform the output using these graphs as oracles for conditional independencies.
-Furthermore, we directly compute the BIC for the outputs, as all PAG members are asymptotically score-equivalent.
+We use three variants of the well-known Sachs dataset (Sachs et al.,, 2005), namely Sachs-5 with 5 variables, Sachs-7 with 7 variables, and the original Sachs with 11 variables, to analyze CD under latent confounding.
+In each variant, the variables with the most false observed independencies were made latent, ensuring that the resulting graphs remain sufficiently connected.
+Further details are in Section C.3.
+Setup.
+We again apply our active elicitation strategy as described in Section 4 with few adjustments.
+A complete description of the LLM elicitation framework can be found in Section D.1.
+We prompt the LLM using simple strategies similar to  Kiciman et al., (2024); Bazaluk et al., (2025).
+For these experiments, each GPT-4o (OpenAI et al.,, 2024) answer is assigned a distinct reliability $\pi$.
+There are many works which discuss about deriving uncertainty over the response of an LLM (Xiong et al.,, 2024; Tanneru et al.,, 2024; Ma et al.,, 2025).
+In this work, we use a Monte Carlo estimate with 10 sampled GPT-4o generations for each query (Bazaluk et al.,, 2025); see Appendix D.
 Results.
-Table 1 compares AGFN against baseline CD algorithms.
-Notably, our method consistently outperforms the only probabilistic baseline in the literature (N-ADMG) in terms of both SHD and BIC.
-As expected, however, the average BIC and SHD induced by AGFN are larger than those induced by the bootstrapped versions of the non-probabilistic algorithms, and the variances are greater; this is due to the inherent sampling diversity of our method and the resulting generation of possibly implausible samples.
-Indeed, Table 2 shows that the three most rewarding samples from AGFN are as good as (and sometimes better than) the other CD algorithms.
-Results for N-ADMG comprise the three most frequent samples from the variational distribution.
-chain4 IV collfork
-FCI $2.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.00}}$ 2.00 $3.83{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.90}}$ 2.90 $5.43{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.87}}$ 1.87
-GFCI $\mathbf{1.50}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.63}}$ 1.63 $3.63{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 3.16}}$ 3.16 $5.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.11}}$ 2.11
-DCD $2.27{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.46}}$ 1.46 $4.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.17}}$ 2.17 $5.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\pm\scriptstyle 2.13}}$ 2.13
-N-ADMG (top 3) $4.38{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 0.81}}$ 0.81 $6.08{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.77}}$ 1.77 $6.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 0.93}}$ 0.93
-AGFN (top 3) $2.00{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.55}}$ 1.55 $\mathbf{3.50}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 3.29}}$ 3.29 $\mathbf{4.90}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.70}}$ 2.70
-Table 2: SHD for point estimates. The mean SHD of the top-3 AGFN draws is comparable to or better than baselines.
-Figure 5: CD with simulated human feedback.
-The top/bottom row shows the mean SHD/BIC of AGFN samples as a function of human interactions.
-Probing the expert about the edge that minimizes the mean cross-entropy leads to a faster decrease in BIC compared to a random strategy.
-The SHD decreases similarly in both cases.
-Results reflect the outcomes of 30 simulations.
-5.3 Simulating humans in the loop
-Data.
-We follow the procedure from Section 5.1 to generate graphs with $4$, $6$, $8$ and $10$ nodes.
-We draw $500$ samples from a compatible linear Gaussian SCM and use them to train an AGFN.
-Then, we follow our active elicitation strategy from Section 4 to probe simulated humans, adhering to the generative model described in the same section, with $\pi=0.9$.
-Setup.
-Since we are the first to propose an optimal design for expert knowledge elicitation, there are no baselines to compare AGFN against.
-That being said, we aim to determine whether the inclusion of expert feedback enhances the concentration of the learned distribution around the true AG, and evaluate the effectiveness of our elicitation strategy.
-To do so, we measure SHD to the true AG and BIC as a function of the number of expert interactions.
-Results.
-Figure 5 shows that incorporating expert feedback substantially decreases the expected SHD and BIC under our belief over AGs.
-On the one hand, the remarkable decrease in expected SHD shows that our belief becomes increasingly focused on the true AG as we iteratively request the expert’s feedback, regardless of the querying strategy.
-On the other hand, the second row shows that our querying strategy results in a substantial decrease in the BIC, demonstrating a faster reduction than random queries.
-This validates the notion that some edges are more informative than others, and we should prioritize them when probing the expert.
-6 Related Work
-CD under latent confounding.
-Following the seminal works by Spirtes et al.
-(2001) and Zhang (2008b) introducing the complete FCI, a variety of works have emerged.
-Among them are algorithms designed for sparse scenarios, including RFCI (Colombo et al., 2012) and others (Silva, 2013; Claassen et al., 2013).
-Notably, Silva (2013)’s framework uses a Bayesian approach to CD of Gaussian causal diagrams based on sparse covariance matrices.
-Nonetheless, it requires sampling one edge at a time and relies on numerical heuristics that might effectively alter the posterior we are sampling from.
-Colombo et al.
-(2012) introduced the conservative FCI to handle conflicts arising from statistical errors in scenarios with limited data, even though it yields less informative results.
-Subsequent efforts to improve reliability led to the emergence of constraint-based CD algorithms based in Boolean satisfiability (Hyttinen et al., 2014; Magliacane et al., 2016), although they are known to scale poorly on $|\mathbf{V}|$ (Lu et al., 2021).
-In another paradigm, score-based search algorithms rank MAGs according to goodness-of-fit measures, commonly using BIC for linear Gaussian SCMs (Triantafillou and Tsamardinos, 2016; Zhalama et al., 2017a; Rantanen et al., 2021).
-There are also hybrid approaches that combine constraint-based strategies to reduce the search space, such as GFCI (Ogarrio et al., 2016), M3HC (Tsirlis et al., 2018), BCCD (Claassen and Heskes, 2012), and GSPo (Bernstein et al., 2020).
-Continuous optimization approaches have recently emerged as a novel approach to score-based CD, such as DCD (Bhattacharya et al., 2021) and N-ADMG Ashman et al.
-(2023).
-CD with expert knowledge.
-Previous works on CD have explored various forms of background knowledge.
-This includes knowledge on edge existence/non-existence Meek (1995b), ancestral constraints Chen et al.
-(2016), variable grouping Parviainen and Kaski (2017), partial order Andrews (2020) and typing of variables (Brouillard et al., 2022).
-Incorporating expert knowledge is pivotal to reducing the search space and the size of the learned equivalence class.
-However, due to significant challenges, up to date, there are only a few works trying to integrate human knowledge into CD within the context of latent confounding (Andrews, 2020; Wang et al., 2022).
-These works operate under the assumption of perfect expert feedback.
-In contrast, our contribution is novel in that it confronts the challenges of real-world situations where expert input might be inaccurate.
+Figure 5 shows that incorporating GPT-4o as an expert leads to a consistent reduction in expected SHD as more feedback is provided.
+Although the LLM is not a perfect expert, occasionally providing incorrect responses, the results exhibit robust performance.
+This highlights that explicitly accounting for varying levels of uncertainty across different LLM responses can substantially improve the accuracy and reliability of our model, even when expert feedback is imperfect.
 6 DISCUSSION
-We presented AGFN, the first probabilistic CD method that accounts for latent confounding and incorporates potentially noisy human feedback in the loop.
+We presented AGFN, the first probabilistic CD method that accounts for latent confounding and incorporates potentially noisy expert feedback in the loop.
 AGFN samples AGs according to a score function, quantifying the uncertainty in the learning process.
-Furthermore, it can leverage human feedback in an optimal design strategy, efficiently reducing our uncertainty on the true data-generating model.
-This work is focused on linear Gaussian models, using BIC as our score.
-However, the implementation of AGFNs is not restricted by this choice.
-In principle, we could replace the BIC with alternative score functions that are more appropriate for different types of variables, e.g.
-for discrete data (Drton and Richardson, 2008).
-It is also important to highlight that our framework does not require retraining the AGFN after we see human feedback.
-Moreover, AGFN is a GPU-powered algorithm, and while we used only one GPU in our experiments, it is possible to greatly accelerate AGFN by using cluster architectures with multiple GPUs.
-By offering uncertainty-quantified CD together with a recipe for including humans in the loop, we expect AGFNs will significantly enhance the accuracy and reliability of CD, especially in real-world domains.
-Moreover, AGFNs bring a novel perspective to developing more comprehensive tools for downstream causal tasks Bareinboim and Pearl (2016), as the resulting distribution encodes knowledge from data and human feedback while accounting for epistemic uncertainty.
-For example, methods for causal reasoning that currently rely on a single AG Zhang (2008a); Jaber et al.
-(2022) could exploit this distribution to incorporate a richer understanding of uncertainty and knowledge, thereby enhancing their robustness and reliability.
-Acknowledgments
-Diego Mesquita acknowledges the support by the Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro FAPERJ (SEI-260003/000709/2023), the São Paulo Research Foundation FAPESP (2023/00815-6), the Conselho Nacional de Desenvolvimento Científico e Tecnológico CNPq (404336/2023-0), and the Silicon Valley Community Foundation through the University Blockchain Research Initiative (Grant #2022-199610).
-António Góis acknowledges the support by Samsung Electronics Co., Ldt.
-Adèle Ribeiro and Dominik Heider were supported by the LOEWE program of the State of Hesse (Germany) in the Diffusible Signals research cluster and by the German Federal Ministry of Education and Research (BMBF) [031L0267A] (Deep Insight).
+Furthermore, it can leverage expert feedback in an optimal design strategy, efficiently reducing our uncertainty on the true data-generating model.
+Notably, AGFNs are not limited to a specific choice of score.
+In principle, the BIC could be replaced with alternative scoring functions better suited for different types of variables, such as discrete data (Drton and Richardson,, 2008).
+Furthermore, our framework does not require retraining the AGFN after incorporating expert feedback, allowing efficient updates.
+By combining uncertainty-quantified CD with a systematic approach for incorporating experts in the loop, AGFNs are expected to substantially improve the accuracy and reliability of CD, particularly in real-world domains.
+Our results further demonstrate that, even when using imperfect experts such as an LLM, the framework successfully converges to the true AG.
+Beyond this, AGFNs offer a novel perspective for developing more comprehensive tools for downstream causal tasks (Bareinboim and Pearl,, 2016), as the resulting distribution integrates both data-driven and expert-informed knowledge while explicitly accounting for epistemic uncertainty.
+For instance, causal reasoning methods that traditionally rely on a single AG (Zhang, 2008a, ; Jaber et al.,, 2022) could leverage this distribution to incorporate richer uncertainty estimates and expert knowledge, thereby enhancing robustness and reliability.
+Acknowledgements
+Tiago da Silva, Eliezer Silva and Diego Mesquita acknowledge the support of Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro FAPERJ (SEI-260003/000709/2023), São Paulo Research Foundation FAPESP (2023/00815-6), Conselho Nacional de Desenvolvimento Científico e Tecnológico CNPQ (404336/2023-0, 445170/2024-7, 305692/2025-9).
+Bruna Bazaluk acknowledges the support of Coordenação de Aperfeiçoamento de Pessoal de Nível Superior – Brasil (CAPES) – Código de Financiamento 001.
+Adèle Ribeiro and Dominik Heider were supported by the LOEWE program of the State of Hesse (Germany) in the Diffusible Signals research cluster and by the German Federal Ministry of Research, Technology and Space (BMBTR) [031L0267A] (Deep Insight).
+António Góis acknowledges the Canada CIFAR AI Chair program (Mila), by an unrestricted gift from Google, and by Samsung Electronics Co., Ldt.
 Samuel Kaski was supported by the Academy of Finland (Flagship programme: Finnish Center for Artificial Intelligence FCAI), EU Horizon 2020 (European Network of AI Excellence Centres ELISE, grant agreement 951847), UKRI Turing AI World-Leading Researcher Fellowship (EP/W002973/1).
-We also acknowledge the computational resources provided by the Aalto Science-IT Project from Computer Science IT.
-Appendix A Additional Related Works
-Generative Flow Networks (GFlowNets; Bengio et al., 2021a, b) are generative models that sample discrete composite objects from an unnormalized reward function.
-They have been successfully used to sample various structures such as protein sequences (Jain et al., 2022) and schedules (Zhang et al., 2023).
-They have also been used to train energy-based models (Zhang et al., 2022).
-In the field of structure learning, they have been applied to Bayesian networks — more specifically to sample a posterior over DAGs in linear Gaussian networks, although without accounting for unobserved confounding (Deleu et al., 2022).
-Recently, Deleu et al.
-(2023) proposed an extension to jointly infer the structure and parameters, also grounded in the assumption of causal sufficiency.
-It is worth highlighting that training GFlowNets in these scenarios presents optimization challenges, resulting in the utilization of a variety of loss functions (Shen et al., 2023).
-Moreover, Lahlou et al.
-(2023) proposed an extension of GFlowNets to continuous domains.
-Appendix B Cross-entropy acquisition
-The expected mutual information and the information gain are the most widely used information-theoretic measures to actively interact with a human and choose the most informative data points to be labeled Ryan et al.
-(2015).
+We also acknowledge the computational resources provided i) by the Aalto Science-IT Project from Computer Science IT, and ii) FGV TIC.
+Appendix A Cross-entropy acquisition
+The expected mutual information and the information gain are the most widely used information-theoretic measures to actively interact with a human and choose the most informative data points to be labeled (Ryan et al.,, 2015).
 However, we instead use the negative expected cross-entropy between the current and updated beliefs as the acquisition function of our experimental design (see eq.
-17).
+13).
 As we show next, the approximation of both the mutual information and the information gain is intrinsically dependent upon the estimation of the log-partition of the updated beliefs over the space of ancestral graphs.
-Doing so is computationally intensive, and we would either need to use a Monte Carlo estimator of the integrals or use some posterior approximation — in both cases, leading to asymptotically biased estimates of the acquisition.
+Doing so is computationally intensive, and we would either need to use a Monte Carlo estimator of the integrals or use some posterior approximation – in both cases, leading to asymptotically biased estimates of the acquisition.
 In contrast, we can easily leverage AGFN samples to compute asymptotically unbiased estimates of our acquisition function.
 The next paragraphs provide further details.
-Mutual information.
+A.1 Mutual information
 The mutual information between two random variables $X$ and $Y$ with joint distribution $p(X,Y)$ and marginal distributions $p(X)$ and $p(Y)$ is
@@
 Thus, the selection criterion in eq.
-20 requires an accurate estimate of $\log c_{K+1}(f_{r})$ — which is well-known for being a difficult problem Ma et al.
-(2013) — and the Monte Carlo estimator for the log-partition function is asymptotically biased.
-Information gain.
+16 requires an accurate estimate of $\log c_{K+1}(f_{r})$, which is well-known for being a difficult problem (Ma et al.,, 2013), and the Monte Carlo estimator for the log-partition function is asymptotically biased.
+A.2 Information gain
 The expected information gain of an elicitation is defined as the expected KL divergence between our updated and current beliefs over ancestral graphs.
-This approach is widely employed in Bayesian experimental design Ryan et al.
-(2015).
+This approach is widely employed in Bayesian experimental design (Ryan et al.,, 2015).
 In our framework, the information gain resulting from a feedback $f_{r}$ is
@@
 20 suffers from the same problems of eq.
-20: it requires approximating the logarithm of the partition function $c_{K+1}(f_{r})$ of a distribution over the combinatorially large space of ancestral graphs, which is notably very challenging to estimate.
+16: it requires approximating the logarithm of the partition function $c_{K+1}(f_{r})$ of a distribution over the combinatorially large space of ancestral graphs, which is notably very challenging to estimate.
 Indeed, as
@@
 $=\underset{\mathcal{G}\sim q(\cdot|\mathbf{f}_{K+1})}{\mathbb{E}}\left[\log p(f_{r}|\omega_{r})+\log c_{K+1}(f_{r})-\log c_{K}\right],$
-with $f_{r_{K+1}}=f_{r}$, $c_{K}$ as the partition function of $q(\cdot|\mathbf{f}_{K})$ — that does not depend upon $f_{r}$ —, and $c_{K+1}(f_{r})$ defined in eq.
-22, the estimation of the information gain is inherently dependent upon the estimation of the log-partition function.
-Cross-entropy.
+with $f_{r_{K+1}}=f_{r}$, $c_{K}$ as the partition function of $q(\cdot|\mathbf{f}_{K})$ that does not depend upon $f_{r}$, and $c_{K+1}(f_{r})$ defined in eq.
+18, the estimation of the information gain is inherently dependent upon the estimation of the log-partition function.
+A.3 Cross-entropy
 The cross-entropy between our updated and current beliefs is an intuitively plausible and practically useful strategy to interact with an expert efficiently. In fact, since
-$\begin{split}\mathbf{H}&[q(\cdot|\mathbf{f}_{K+1}),q(\cdot|\mathbf{f}_{K})]\\ &=\underset{\mathcal{G}\sim q(\cdot|\mathbf{f}_{K+1})}{\mathbb{E}}[-\log q(\mathcal{G}|\mathbf{f}_{K})]\\ &=\underset{\mathcal{G}\sim q(\cdot|\mathbf{f}_{K+1})}{\mathbb{E}}\left[-\log p_{\theta}(\mathcal{G})-\sum_{1\leq k\leq K}\log p(\omega_{r_{k}}|f_{r_{k}})-l_{K}\right]\end{split},$
-in which $l_{K}$ is the log-partition function of the distribution $q(\cdot|\mathbf{f}_{K})$.
+$\begin{split}\mathbf{H}&[q(\cdot|\mathbf{f}_{K+1}),q(\cdot|\mathbf{f}_{K})]\\ &=\underset{\mathcal{G}\sim q(\cdot|\mathbf{f}_{K+1})}{\mathbb{E}}[-\log q(\mathcal{G}|\mathbf{f}_{K})]\\ &=\underset{\mathcal{G}\sim q(\cdot|\mathbf{f}_{K+1})}{\mathbb{E}}\left[-\log p_{\theta}(\mathcal{G})-\sum_{1\leq k\leq K}\log p(\omega_{r_{k}}|f_{r_{k}})-l_{K}\right]\end{split},$ (21)
+in which $l_{K}$ is the log-partition function of the distribution $q(\cdot|\mathbf{f}_{K})$, there is no need to estimate any partition function for selecting a query with which to probe the expert.
 Further, the cross-entropy depends exclusively upon i) the logarithm of the samples’ rewards, $\log p_{\theta}(\mathcal{G})$, which is readily computed within AGFN’s generative process, and ii) the posterior distribution over the relations’ features $\omega_{r}$ given the expert’s feedbacks $f_{r}$, which is available in closed form.
-Hence, the previously mentioned expectation is unbiasedly and consistently estimated by our importance sampling scheme.
+Hence, the previously mentioned expectation in Equation 21 is unbiasedly and consistently estimated through a Monte Carlo estimator based on samples drawn from AGFN’s updated distribution.
 Furthermore, our empirical findings in fig.
 12 suggest that the cross-entropy yields good results and consistently outperforms a uniformly random strategy with respect to the BIC score.
+Appendix B Algorithmic design of AGFN
+To keep track of the invalid actions leading to non-ancestral graphs within the generative process, we follow an approach similar to Giudici and Castelo, (2003); Deleu et al., (2022) and decompose the mask $\mathbf{m}_{t}$ of valid actions at the $t$th state $G_{t}$ into a term corresponding to the state’s adjacency matrix and a term resembling the state’s transitive closure.
+Formally, we define
+  1. 1.
+    $\alpha_{t}\colon\mathbf{V}\times\mathbf{V}\rightarrow\{1,0\}$ as an indicator function verifying whether there is a directed path between the nodes $v_{i}\in V$ and $v_{j}\in\mathbf{V}$ in the transpose graph $\mathcal{G}_{t}^{\intercal}$; and
+  2. 2.
+    $\beta_{t}\colon\mathbf{V}\times\mathbf{V}\rightarrow\{1,0\}$ as another indicator function identifying whether there is, within the transpose graph, a path between $v_{i}\in\mathbf{V}$ and $v_{j}\in\mathbf{V}$ with all, except one, directed edges pointing to $v_{j}$.
+In this context, we note that adding a directed edge $v_{i}\rightarrow v_{j}$ leads to an cycle if and only if $\alpha_{t}(v_{i},v_{j})=1$, whereas it leads to an almost cycle if and only if $\beta_{t}(v_{i},v_{j})=1$.
+Correlatively, the addition of a bidirected edge $v_{i}\leftrightarrow v_{j}$ induces an almost cycle within the current state if and only if either $\alpha_{t}(v_{i},v_{j})=1$ or $\alpha_{t}(v_{j},v_{i})=1$.
+In this context, as an action is allowed if and only if it doesn’t lead to the formation of a directed or an almost directed cycle, we split the mask $\mathbf{m}_{t}=\mathbf{d}_{t}\oplus\mathbf{b}_{t}$ as a concatenation into actions corresponding to adding directed edges $\mathbf{d}_{t}$ and actions corresponding to bidirected edges $\mathbf{b}_{t}$, decomposing it as
+$\begin{split}&\mathbf{d}_{t}=\mathbf{E}_{t}\odot\left(1-\boldsymbol{\alpha}_{t}\right)\odot\left(1-\boldsymbol{\beta}_{t}\right)\\ &\mathbf{b}_{t}=\mathrm{triu}\left(\mathbf{E}_{t}\odot\left(1-\boldsymbol{\alpha}_{t}\right)\odot\left(1-\boldsymbol{\alpha}_{t}^{\intercal}\right)\right)\end{split},$ (22)
+with $\mathbf{E}_{t}=\left(1-\mathbf{D}_{t}\right)\odot\left(1-\mathbf{B}_{t}\right)$ denoting the unconnectedness of each pair of nodes; $\boldsymbol{\alpha}_{t}$ (resp.
+$\boldsymbol{\beta}_{t})$) representing a matrix with entries $\alpha_{t}(v_{i},v_{j})$ (resp.
+$\beta_{t}(v_{i},v_{j})$); and $\mathrm{triu}$ corresponding to the operation of extracting the upper triangular part of a matrix.
+Notably, each of the terms of the preceding equation can be efficiently updated as we iteratively build the ancestral graphs.
+For $\mathbf{E}_{t}$, we simply update the entries corresponding to the chosen action.
+For $\boldsymbol{\alpha}_{t}$, notice that adding an edge $v_{i}\rightarrow v_{j}$ creates a directed path from the node $w_{j}$ to the node $w_{i}$ in $\mathcal{G}_{t}^{\intercal}$ if and only if there is a directed path from $w_{j}$ to $v_{j}$ and a directed path from $v_{i}$ to $w_{i}$ in $\mathcal{G}_{t}^{\intercal}$, with a similar correspondence for directed paths with single bidirected edges; thus, after adding $w_{i}\rightarrow w_{j}$, we update
+$\begin{split}&\boldsymbol{\alpha}_{t+1}=\boldsymbol{\alpha}_{t}+\boldsymbol{\alpha}_{t}(\cdot,v_{j})\boldsymbol{\alpha}_{t}(v_{i},\cdot)\\ &\boldsymbol{\beta}_{t+1}=7\boldsymbol{\beta}_{t}+\boldsymbol{\beta}_{t}(\cdot,v_{j})\boldsymbol{\alpha}_{t}(v_{i},\cdot)+\boldsymbol{\alpha}_{t}(\cdot,v_{j})\boldsymbol{\beta}_{t}(v_{i},\cdot)\end{split},$ (23)
+in which $\boldsymbol{\alpha}_{t}(\cdot,v)$ (resp.
+$\boldsymbol{\alpha}_{t}(v,\cdot)$) corresponds to the column (resp.
+row) vector indexed by $v$ in $\boldsymbol{\alpha}_{t}$ and the operator ++ represents a bitwise or.
+By a similar reasoning, the addition of a bidirected edge $v_{i}\leftrightarrow v_{j}$ corresponds to the update rule
+$\begin{split}&\boldsymbol{\alpha}_{t+1}=\boldsymbol{\alpha}_{t}\\ &\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\boldsymbol{\alpha}_{t}(\cdot,v_{i})\boldsymbol{\alpha}_{t}(v_{j},\cdot)+\boldsymbol{\alpha}_{t}(\cdot,v_{j})\boldsymbol{\alpha}_{t}(v_{i},\cdot)\end{split}.$ (24)
+Alternatively, albeit significantly more costly, one could enumerate the potential children of a state and use Bhattacharya et al., (2021)’s algebraic constraint to decide the validity of the corresponding transitions.
 Appendix C Experimental details
-We lay out the experimental and implementational details of our empirical analysis in the next subsections.
-In Section C.1, we describe the specific configurations of the CD algorithms that we compared with our method in table 2.
-Then, we consider in section C.2 some practical guidelines and architectural specifications that enable us to train and make inferences with AGFN efficiently.
-Finally, we contemplate in section C.3 the algorithmic details for simulating the expert’s feedback according to our model for active knowledge elicitation.
+We lay out more experimental and implementation details of our empirical analysis in the next subsections.
 C.1 Baselines
-FCI.
-For the results in table 1, we first estimated a PAG using the stable version of FCI, which produces a fully order-independent final skeleton (Colombo et al., 2014).
+C.1.1 FCI
+We first estimated a PAG using the stable version of FCI, which produces a fully order-independent final skeleton (Colombo et al.,, 2014).
 To identify conditional independencies, we used Fisher’s Z partial correlation test with a significance level of $\alpha=0.05$.
 The BIC score associated with the PAG estimated by the FCI was computed as the BIC of a randomly selected maximal AG (MAG) within the equivalence class characterized by such PAG.
-The maximality of an AG depends on the absence of inducing paths between non-adjacent variables, which are paths where every node along it (except the endpoints) is a collider and every collider is an ancestor of an endpoint (Rantanen et al., 2021).
+The maximality of an AG depends on the absence of inducing paths between non-adjacent variables, which are paths where every node along it (except the endpoints) is a collider and every collider is an ancestor of an endpoint (Rantanen et al.,, 2021).
 This ensures that in the MAG every non-adjacent pair of nodes is m-separated by some set of other variables.
-Importantly, Markov equivalent MAGs exhibit asymptotic equivalence in terms of BIC scores (Richardson and Spirtes, 2002).
+Importantly, Markov equivalent MAGs exhibit asymptotic equivalence in terms of BIC scores (Richardson and Spirtes,, 2002).
 As a result, the choice of a random MAG does not disrupt the validity of our results.
-GFCI.
+C.1.2 GFCI
 Similarly, we applied GFCI with an initial search algorithm (FGS) based on the BIC score and the subsequent application of the FCI with conditional independencies identified by the Fisher’s Z partial correlation test with a significance level $\alpha=0.05$.
-This was performed for all datasets listed in table 1.
 Also similar to the procedure adopted with the FCI, the BIC score associated with the estimated PAG was computed as the BIC of a randomly selected MAG within the equivalence class characterized by such PAG.
-DCD.
+C.1.3 ACI
+We used the implementation provided by the authors at GitHub [^1] ACI repository, available online at https://github.com/caus-am/aci.
+For the Answer Set Programming solver, we used clingo 4, version 5.6.2, also available at GitHub  [^2] Clingo repository, available online at https://github.com/potassco/clingo.
+Similarly to the FCI, we tested conditional indepedencies using Fisher’s Z partial correlation test with a significance level of $\alpha=0.05$.
+ACI also outputs a PAG, so we computed the BIC of a randomly selected MAG within the represented Markov equivalence class.
+C.1.4 DCD
 We adhered to the instructions provided in the official repository [^3] Available online at https://gitlab.com/rbhatta8/dcd.
-to apply the DCD method on the datasets in table 1.
+to apply the DCD method.
 The SHD was obtained between the ground-truth PAG and the PAG corresponding to the estimated ADMG (i.e., the one obtained via FCI by using the d-separations entailed by the estimated ADMG as an oracle for conditional independencies).
 On the other hand, the BIC was computed for the estimated ADMG directly.
-N-ADMG.
+C.1.5 N-ADMG
 To estimate the parameters of the variational distribution defined by N-ADMG, we executed the code provided at the official repository [^4] Available online at https://github.com/microsoft/causica/releases/tag/v0.0.0.
-For fairness, we used the same hyperparameters and architectures reported in their original work Li et al.
-(2023); in particular, we trained the models for $30$k epochs.
-After this, we sampled $100$k graphs from the learned distribution.
+For fairness, we used the same hyperparameters and architectures reported in their original work (Li et al.,, 2023); in particular, we trained the models for 3030k epochs.
+After this, we sampled 100100k graphs from the learned distribution.
 It is worth mentioning that the constraints of bow-free ADMG are guaranteed in the N-ADMG samples only in an asymptotic sense.
@@
 Then, we proceeded exactly as with DCD to estimate both the average SHD and the average BIC under the variational distribution.
-C.2 Implementational details for AGFN
+C.1.6 DAG-GFlowNets
+DAG-GFlowNets (Deleu et al.,, 2022) are a GFlowNet-based method for structure learning of Bayesian networks.
+In a nutshell, this family of models first defines a generative process of DAGs by iteratively adding edges to an initially edgeless graph.
+Then, a GFlowNet is trained to sample states proportionally to a goodness-of-fit score function (Deleu et al.,, 2023).
+Importantly, when regarded as a CD algorithm, DAG-GFlowNets ignore latent confounding.
+To illustrate this limitation, we consider three distinct causal graphs belonging to MEC with invariant bidirected edges (see diagrams 1, 2, and 3 in Figure 6).
+Under these conditions, we show that AGFNs are often more accurate than DAG-GFlowNets.
+Results To evaluate the score function in Equation 27, we considered datasets with 500 samples.
+Table 2 below compares the SHD of point estimates for DAG-GFlowNets with AGFN’s.
+Notably, AGFN outperforms DAG-GFlowNet in all cases.
+This is expected, since DAG-GFlowNets, in contrast with AGFNs, are incapable of sampling/learning graphs with bidirected edges, even under faithfulness.
+DAG-GFlowNet AGFN
+Causal diagram 1 $3.80$ $\mathbf{1.67}$
+Causal diagram 2 $6.40$ $\mathbf{5.27}$
+Causal diagram 3 $8.22$ $\mathbf{7.87}$
+Table 2: AGFN outperforms DAG-GFN under the presence of latent confounders. See Figure 6 for the diagrams.
+(a) CD 1.
+(b) CD 2.
+(c) CD 3.
+Figure 6: Causal diagrams under latent confounding for comparing AGFN with DAG-GFlowNet. See results in Table 2.
+C.2 Implementation details for AGFN
+C.2.1 Forward flow
+We use a Graph Isomorphism Network (Xu et al.,, 2019) $\Phi$ to compute a $d$-dimensional representation for each node in the AG $\mathcal{G}_{t}$ at the $t$-th step of the generative process and use sum pooling to get an embedding for $\mathcal{G}_{t}$.
+Let $\mathcal{A}_{t}$ denote the space of feasible actions at $\mathcal{G}_{t}$ (i.e., those leading to valid AGs).
+We then map $\mathcal{G}_{t}$’s embedding to a distribution over $\mathcal{A}_{t}$ using a multilayer perceptron (MLP) $\phi\colon\mathbb{R}^{d}\rightarrow\mathbb{R}^{|\mathcal{A}_{t}|}$ with a softmax activation at its last layer.
+Specifically, given $\mathbf{H}^{(t)}=\Phi(\mathcal{G}_{t})\in\mathbb{R}^{|\mathbf{V}|\times d}$, we compute
+$\mathbf{p}=\phi\left(\sum\nolimits_{v\in\mathbf{V}}\mathbf{H}_{v}^{(t)}\right)$ (25)
+as the probability distribution over the actions at $\mathcal{G}_{t}$.
+C.2.2 Backward flow
+Backward actions correspond to removing edges.
+Following Shen et al., (2023), we parametrize the backward flow $F_{B}$ with an MLP and alternate between updating $\pi_{F}$ and $\pi_{F_{B}}$, using gradient-based optimization.
 Figure 7: Tempered rewards.
 Training AGFN to sample from increasingly cold distributions (eq.
-25) enables us to increase the proportion of high-scoring graphs (i.e., with a low BIC-score) with the drawback of reducing the AGFN’s sampling diversity.
-Masking.
+28) enables us to increase the proportion of high-scoring graphs (i.e., with a low BIC-score) with the drawback of reducing the AGFN’s sampling diversity.
+C.2.3 Masking
 To ensure AGFN only samples ancestral graphs, we keep track of a binary mask $\mathbf{m}_{t}$ that indicates which actions lead to a valid state at the iteration $t$ of the generative process; this mask defines the support of the policy evaluated at the corresponding state.
-In more detail, let $\mathbf{y}_{t}$ be the last layer embedding (prior to a softmax) at iteration $t$ of the neural network used to parametrize the forward flow of AGFN.
+See Appendix B for further details regarding the computation of this mask.
+In practice, we let $\mathbf{y}_{t}$ be the last layer embedding (prior to a softmax) at iteration $t$ of the neural network used to parametrize the forward policy of AGFN.
 The probability distribution over the space of feasible actions is then
 $\mathbf{p}_{t}=\text{Softmax}\left(\mathbf{y}_{t}\odot\mathbf{m}_{t}+\epsilon\cdot(1-\mathbf{m}_{t})\right)$
-for a large and negative constant $\epsilon$. We empirically verified that $\epsilon=-10^{5}$ is sufficient to avoid the sampling of non-ancestral graphs.
-Exploratory policy.
+for a large and negative constant $\epsilon$.
+We empirically verified that $\epsilon=-10^{5}$ is sufficient to avoid the sampling of non-ancestral graphs when using a double-precision floating point format.
+C.2.4 Exploratory policy
 During training, we must use an exploratory policy that (i) enables the exploration of yet unvisited states within the pointed DAG and (ii) exploits highly valuable and already visited states.
-To alleviate this phenomenon, we also draw trajectories from a uniform policy, which is a widespread practice in the literature (Bengio et al., 2021a; Deleu et al., 2022; Shen et al., 2023).
+To satisfy these criteria, we also draw trajectories from a uniform policy, which is a widespread practice in the literature (Bengio et al., 2021a, ; Deleu et al.,, 2022; Shen et al.,, 2023).
 More precisely, let $\text{Ch}(\mathcal{G}_{t})$ be the set of states (i.e., ancestral graphs) directly reachable from $\mathcal{G}_{t}$ and $\alpha\in[0,1]$.
@@
 During inference, we set $\alpha=1$ to sample actions exclusively from the GFlowNet’s learned policy.
-Detection of invalid states.
-We use the algebraic condition in eq.
-8 to check whether a graph $\mathcal{G}$ is ancestral.
-At each iteration of the generative process, we draw an action from the current exploratory policy and test the ancestrality of the updated graph; if it is not ancestral, we revert the sampled action and mask it.
-Importantly, this protocol guarantees that all graphs sampled from AGFN are ancestral.
-Batch sampling.
+C.2.5 Validating AGFN’s samples
+To validate that AGFN’s sampled graphs are ancestral, we use the algebraic condition proposed by Bhattacharya et al., (2021), namely,
+$\textrm{trace}(e^{D})-d+\mathbf{1}^{T}e^{D\odot B}\cdot\mathbf{1}=0,$ (26)
+in which $D\in\{1,0\}^{d}$ is the graph’s adjacency matrix corresponding to the directed edges and $B$ is the one corresponding to the bidirected edges; $\mathbf{1}$ is a vector of ones; and $\odot$ represents Hadamard’s pointwise product.
+C.2.6 Batch sampling
 We exploit batch sampling to fully leverage the power of GPU-based computing in AGFN.
 As both the maximum-log-likelihood-based reward and the validation of the states are parallelizable operations, we are able to distribute them across multiple processing units and efficiently draw samples from the learned distribution.
-Crucially, this end-to-end parallelization substantially improves the computational feasibility of our algorithm and is a notable feature generally unavailable in prior works Zhang (2008b); Ogarrio et al.
-(2016); Rantanen et al.
-(2021).
-We use a batch size of 256 for all the experiments — independently of the graph size.
+Crucially, this end-to-end parallelization substantially improves the computational feasibility of our algorithm and is a notable feature generally unavailable in prior works (Zhang, 2008b, ; Ogarrio et al.,, 2016; Rantanen et al.,, 2021).
+We use a batch size of 128 trajectories for all the experiments, independently of the graph size.
 Figure 8: Sensitivity of our active knowledge elicitation framework to the reliability of the expert.
@@
 As expected, the improvements entailed by the expert’s feedback become increasingly effective as we increase the expert’s reliability from $0.1$ to $0.9$.
-Results reflect the outcome of 30 scenarios simulated accordingly to algorithm 1 with a random canonical diagram $\mathcal{G}^{\star}$ with $5$ nodes.
+Results reflect the outcome of 30 scenarios simulated accordingly to algorithm 1 with a random canonical diagram $\mathcal{G}^{\star}$ with 55 nodes.
 We used our active knowledge elicitation scheme to select the query at each iteration.
-Training hyperparameters.
-For AGFN’s forward flow, we use a Graph Isomorphism Network (GIN, Xu et al.
-(2019)) with $2$ layers to compute embeddings of dimension $256$.
-Then, we project these embeddings to a probability distribution using a three-layer MLP having leaky RELUs with a negative slope of $-0.01$ as activation functions.
+C.2.7 Training hyperparameters
+For AGFN’s forward flow, we use a Graph Isomorphism Network (GIN, Xu et al., (2019)) with $2$ layers to compute embeddings of dimension $256$.
+Then, we project these embeddings to a probability distribution using a three-layer MLP having leaky RELUs with a negative slope of 0.01-0.01 as activation functions.
 Correspondingly, we use an equally configured three-layer MLP to parametrize AGFN’s backward flow.
@@
 Moreover, we trained the neural networks for $3000$ epochs for the human-in-the-loop simulations (in which we considered graphs having up to $10$ nodes) and for $500$ epochs for both the assessment of the distributional quality of AGFN and the comparison of AGFN with alternative CD approaches.
-Computational settings.
-We trained the AGFNs for the experiments in fig.
-3 and table 1 and fig.
-5 for $500$ epochs in computers equipped with NVIDIA’s V100 GPUs.
-All the experiments were executed in a cluster of NVIDIA’s V100 GPUs and the algorithms were implemented using the machine learning framework PyTorch.
-To estimate the PAG corresponding to AGFN’s samples and compute the SHDs reported in table 1, we used the FCI’s implementation of the pcalg package in R considering the d-separations entailed by these samples as a criterion for conditional dependence.
-C.3 Human in the loop
-Algorithm 1 Simulating humans in the loop
-$\{\mathcal{G}_{t}\}_{1\leq t\leq T}\text{ samples from AGFN}$, $\mathcal{G}^{*}=(\mathbf{V},E)$ true ancestral graph, $\pi$ reliability of the expert’s feedback
-$p(\omega_{r}=k)\leftarrow\frac{1}{T}\sum_{1\leq t\leq T}1_{\{\omega_{r}=k\text{ in }G_{t}\}}\forall k\in[4],r\in{\mathbf{V}\choose 2}$
-$\mathbf{f}\leftarrow\{\}$triangleright Set of feedbacks (answers)
-$\boldsymbol{r}\leftarrow\{\}$triangleright Set of queries (questions)
-$K\leftarrow 1$
-$\omega_{r}^{*}\leftarrow\text{ relation }r\text{'s feature in }\mathcal{G}^{*}\ \forall r\in{\mathbf{V}\choose 2}$
-while $\boldsymbol{r}\neq{\mathbf{V}\choose 2}$ do $\triangleright$ Iteratively request the feedback
-     $r_{K}\leftarrow\underset{r\in{\mathbf{V}\choose 2}\setminus\boldsymbol{r}}{\operatorname*{arg\ max}}\ \underset{{f_{r}\sim p(\cdot)}}{\mathbb{E}}[-\mathbf{H}\left(q(\mathcal{G};\mathbf{f}\cup\{f_{r}\}),q(\mathcal{G};\mathbf{f})\right)]$
-     $\mathbf{r}\leftarrow\mathbf{r}\cup\{r_{K}\}$
-     $f_{K}\sim\text{Cat}\left(\pi\cdot\delta_{\omega_{r_{K}}^{*}}+\left(\frac{1-\pi}{3}\right)\cdot(1-\delta_{\omega_{r_{K}}^{*}})\right)$
-     $\mathbf{f}\leftarrow\mathbf{f}\cup\{f_{K}\}$
-     $K\leftarrow K+1$
-end while
-Algorithmic details.
-We describe in algorithm 1 our procedure for simulating interactions with an expert.
-Initially, we estimate the marginal probabilities $p(\omega_{r}=k)$ of a relation $r$ displaying the feature $k\in\{1,2,3,4\}$ under AGFN’s learned distribution.
-This is our prior distribution.
-In algorithm 1, we denote $\{1,2,3,4\}$ by $[4]$.
-Then, we iteratively select the relation that maximizes our acquisition function; the simulated human thus returns a feedback that equals the selected relation’s true feature with probability $\pi$ or is otherwise uniformly distributed among the incorrect alternatives.
-Importantly, this iterative mechanism can be interrupted at any iteration and the collected feedbacks can be used to compute the importance weights necessary for estimating expectations of functionals under our updated beliefs.
-Appendix D Additional experiments
+C.2.8 Computational settings
+We trained the AGFNs for the experiments for $500$ epochs in computers equipped with NVIDIA’s V100 GPUs.
+All the experiments were executed in a cluster of NVIDIA’s V100 GPUs and the algorithms were implemented using the machine learning framework PyTorch (Paszke et al.,, 2019).
+To estimate the PAG corresponding to AGFN’s samples and compute the SHDs, we used the FCI’s implementation of the pcalg package in R considering the d-separations entailed by these samples as a criterion for conditional dependence.
+C.2.9 Running times for AGFN
+Figure 9 highlights that the required training and sampling times for AGFN scale roughly as a quadratic function of the number of observed variables.
+Correspondingly, the running time for drawing independent graphs from the learned distribution increases linearly in the number of graphs.
+Contrastingly, the cost of SOTA CD algorithms grows exponentially in the number of variables (Zhang, 2008b, ; Ogarrio et al.,, 2016), and sampling from the bootstrapped distribution is thus prohibitively expensive.
+Figure 9: Running times for AGFN.
+The avg.
+time AGFN takes to train/sample scales $\approx$ quadratically in the number of observed variables ($\leftarrow$, center).
+Similarly, the sampling time grows linearly with the number of sampled graphs ($\rightarrow$).
+C.2.10 Score for linear Gaussian models
+For experiments with linear Gaussian models, we use the extended Bayesian Information Criterion (Foygel and Drton,, 2010) as our score function.
+Specifically, for any AG $\mathcal{G}=(\mathbf{V},\mathbf{E})$:
+$U(\mathcal{G})$ $=-2l_{N}(\hat{\mathbf{B}},\hat{\boldsymbol{\Omega}})+|\mathbf{E}|\log n+2|\mathbf{E}|\log|\mathbf{V}|,$ (27)
+[…]
+C.3 Sachs dataset details
+Sachs (Sachs et al.,, 2005) is a dataset that measures the levels of specific proteins and phospholipids in the human cell.
+The data are continuous and the $11$ observed variables/features are: "Raf", "Mek", "Plcg", "PIP2", "PIP3", "Erk", "Akt", "PKA", "PKC", "P38", "Jnk".
+As we would like the dataset to contain latent counfounders, we used three different versions in our experiments: Sachs, Sachs-7 and Sachs-5.
+All have an equal number of samples, $7466$, which is the original amount.
+Sachs is the original dataset.
+In order to reduce the number of variables, we apply a simple conditional independence test and remove the ones with the most false independencies.
+Thus, we ensure the resulting graphs remain sufficiently connected and not overly sparse.
+Sachs-7 contains the following variables: "Mek", "Plcg", "PIP2", "PIP3", "PKA", "PKC", "Jnk" and Sachs-5: "Mek", "PIP3", "PKA", "PKC", "Jnk".
 Figure 10: Architectural design of AGFN.
-Top: The inductively biased parametrization of AGFN’s forward flow — based upon a GNN — enables the substantial reduction of the number of epochs required for training.
+Top: The inductively biased parametrization of AGFN’s forward flow, based upon a GNN, enables the substantial reduction of the number of epochs required for training.
 Bottom: The use of a parametrized backward policy similarly enhances the training efficiency compared to a uniform policy.
 For both experiments, we considered $\mathcal{L}(\theta)<0.1$ as the early stopping criterion to interrupt AGFN’s training.
-Trade-off between diversity and optimality in AGFN.
+C.4 Learning a distribution over sparse AGs
+To illustrate the efficiency and effectiveness of AGFN, we randomly select a 25-node graph and generate a 1000-sized dataset according to the corresponding SCM.
+Subsequently, we modified AGFN’s generative process by constraining the maximum degree a node can have to $3$, which drastically reduces the space of searchable graphs and is easily enforced by modifying the masking procedure described in Appendix B.
+Under these conditions, Figure 11 shows that AGFN accurately learns to sample from the target.
+Figure 11: AGFN learns an accurate approximation to a distribution over 25-node sparse AGs. Left and right axes represent the same quantities at Figure 3’s left-most plot.
+C.4.1 Trade-off between diversity and optimality in AGFN
 We may use tempered rewards to increase the frequency of high-scoring samples and thereby reduce the diversity of AGFN’s distribution. More precisely, we choose a temperature $T$ and consider
 $R_{T}(\mathcal{G})=R(\mathcal{G})^{1/T}=\exp\left\{\frac{\mu-U(\mathcal{G})}{T\sigma}\right\}$ (28)
-as the reward upon which the GFlowNet is trained; if $T\rightarrow 0$, the distribution $p_{T}\propto R_{T}$ converges to a point mass in $R(\mathcal{G})$’s mode and, if $T\rightarrow\infty$, $p_{T}$ converges to an uniform distribution.
-This approach resembles the simulated tempering scheme commonly exploited in Monte Carlo methods Marinari and Parisi (1992) and was previously considered in the context of GFlowNets by Zhang et al.
-(2023).
+as the reward upon which the GFlowNet is trained; if $T\rightarrow 0$, the distribution $p_{T}\propto R_{T}$ converges to an uniform distribution over $R(\mathcal{G})$’s modes and, if $T\rightarrow\infty$, $p_{T}$ converges to an uniform distribution (Geman and Geman,, 1984, Theorem B).
+This approach resembles the simulated tempering scheme commonly exploited in Monte Carlo methods Marinari and Parisi, (1992) and was previously considered in the context of GFlowNets by Zhang et al., (2023).
 Figure 7 shows that progressively cold distributions (i.e., with $T\rightarrow 0$) lead to progressively concentrated and decreasingly diverse samples.
-Notably, the use of cold distributions may be adequate if we are highly confident in our score and are mostly interested in high-scoring samples (e.g., as in Rantanen et al., 2021).
-Figure 9: Human-aided AGFN significantly outperforms alternative CD algorithms.
-Updating AGFN’s distribution according to the feedback of an oracle substantially improves AGFN’s capacity to correctly identify the true ancestral graph; indeed, a single feedback is sufficient to yield results better than (or indistinguishable from) alternative CD algorithms.
-We select the sampled AG with the highest posterior reward as a point estimate of AGFN and use the same datasets listed in table 2.
-The plots summarize the results of $30$ HITL simulations using $\pi=0.9$ and an oracle as an expert (see algorithm 1).
-Sensitivity analysis for different noise levels.
+Notably, the use of cold distributions may be adequate if we are highly confident in our score and are mostly interested in high-scoring samples (e.g., as in Rantanen et al.,, 2021).
+C.5 Sensitivity analysis for different noise levels
 Figure 8 displays the effect of the feedback of an increasingly reliable expert over the expectations of both the SHD and the BIC.
@@
 We emphasize that our model for the expert’s responses is straightforwardly extensible to accommodate multiple feedbacks about the same causal relation under different reliability levels.
-Ablation studies.
+C.6 Ablation studies
 Figure 10 shows the increase in the training efficiency due to our architectural designs for parametrizing both the forward and backward flows of AGFN.
-Noticeably, the use of a two-layer graph isomorphism network (GIN; Xu et al., 2019) with a 256-dimensional embedding for the forward flow entailed a decrease of more than 10x in the number of epochs required for successfully training AGFN; this highlights the effectiveness of an inductively biased architectural design for the parametrization of GFlowNet’s flows.
-Correlatively, the use of a parametrized backward flow significantly enhances the training efficiency of AGFN and emphasizes the inadequacy of a uniformly distributed backward policy pointed out in a previous work Shen et al.
-(2023).
-Human-aided AGFN versus alternative CD methods.
-Figure 9 exposes the significant enhancement of AGFN’s point estimates entailed by our HITL framework for CD.
-This underlines the usefulness of the elicited knowledge, which is simply incorporated into our model through a re-weighting of the reward function, enabling the identification of the true ancestral graph.
-In contrast, most alternative CD algorithms cannot be as easily adapted to include various forms of expert knowledge — and such incorporation, when it is possible, usually precedes any inferential process Andrews (2020) or assumes the knowledge is perfect Wang et al.
-(2022).
+Noticeably, the use of a two-layer graph isomorphism network (GIN; Xu et al.,, 2019) with a 256-dimensional embedding for the forward flow entailed a decrease of more than 10x in the number of epochs required for successfully training AGFN; this highlights the effectiveness of an inductively biased architectural design for the parametrization of GFlowNet’s flows.
+Correlatively, the use of a parametrized backward flow significantly enhances the training efficiency of AGFN and emphasizes the inadequacy of a uniformly distributed backward policy pointed out in a previous work (Shen et al.,, 2023).
+Yeast E. coli
+N-ADMG $1762.50{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 448.51}} 22$$2217.61{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 228.28}}$
+DCD $20941.02{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 127.62}}$ $21994.54{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 138.08}}$
+GFCI $21000.43{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 183.23}}$ $21920.66{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 135.51}}$
+FCI $1250.00{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 438.83}} 22$$2102.09{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 88.98}}$
+ACI $1215.77{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 246.04}} 22$$2071.27{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 188.83}}$
+AGFN $\mathbf{20918.87}{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 71.89}}$ $\mathbf{21852.23}{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 45.29}}$
+Table 3: BIC for GRN datasets. AGFN’s AGs have significantly larger scores (smaller BIC) than baselines’.
+C.7 Experiments on gene regulatory networks (GRN)
+Table 3 shows that AGFN leads to significantly better results in terms of BIC than some of the considered baselines in the structure learning of the 10-node GRNs of a yeast of the E.
+coli bacteria.
+This observation underlines the effectiveness of our method in real-world datasets.
+Moreover, the results in Table 3 attest that AGFN may be successfully scaled beyond the relatively small 5-node causal diagrams considered in the main body of text.
+For this experiment, we implemented a GFlowNet using the same architecture outlined in section C.2 to sample proportionally to the BIC score based on 200 data points.
+The datasets were obtained from the DREAM3 challenge.
+C.8 Assessment of our elicitation framework
+Figure 12 shows that our entropy-minimizing elicitation strategy leads to consistently faster reduction in the BIC when compared against an algorithm that selects the next query uniformly at random.
+In practice, this emphasizes that our approach for probing the expert significantly improves the convergence speed of the expert-in-the-loop process towards better (higher-scoring) AGs.
+In addition, the algorithm we propose incurs a minimal bottleneck to the iterative refinement of a trained AGFN.
+That said, the development of provably optimal experimental designs for probing an expert in the context of CD remains open.
+chain4 IV collfork
+BIC
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 4: Average BIC.
+The (above the dashed line) denotes methods yielding point estimates, for which we employ bootstrap to report the mean and average standard deviation.
+For N-ADMG and AGFN, we estimate the quantities using 100100k sampled graphs.
+In most cases, AGFN incurs higher sample diversity than bootstrapped non-probabilistic CD, and achieves comparable results to these methods in terms of BIC (lower is better).
+Figure 12: CD with simulated human feedback on synthetic graphs.
+Top/bottom row shows the mean SHD/BIC of AGFN samples as a function of human interactions.
+Probing the expert about the edge that minimizes the mean cross-entropy leads to a faster decrease in BIC than a random strategy.
+SHD decreases similarly in both cases.
+Results reflect the outcomes of 30 simulations.
+Appendix D Details on the Experiments with LLMs
+For the preliminary experiments, we used both GPT-4o (OpenAI et al.,, 2024) and Deepseek (V3 and R1) (DeepSeek-AI et al.,, 2024).
+However, Deepseek’s outputs were too far from the correct edges.
+As the feedback was misleading the algorithm, the final results were worse than using AGFN without any feedback.
+As GPT-4o’s feedback was much closer to the real edges, though far from perfect, we decided to report only these results.
+Our LLM-based framework is built using DSPy (Khattab et al.,, 2022, 2024), a library to build modular AI software. We define to the LLM what the input and output will be:
+description=(
+"Pair of nodes about which we want to know the relationship (defined by EdgeType). "
+"Each pair of nodes may be directly associated, via direct causation, or indirectly associated, via latent confounding. "
+"The nature of each node and the underlying problem is described in the context field."
+context: str = dspy.InputField(
+description=(
+"The context of the problem, including the nature of each node (represented by a string) "
+"and the nature of the kind of relationships we are looking for."
+relationship: EdgeType = dspy.OutputField(
+description=(
+"This is the feedback. It describes the nature of the relationship between the input nodes. "
+"This information should be inferred from the provided context, background knowledge, and the nature of the problem."
+Then, for each given query, we prompt the LLM:
+‘You are an expert on the human immune system cell.
+You are investigating the cause-and-effect relationships between a specific set of observed variables representing proteins and phospholipids: "Raf", "Mek", "Plcg", "PIP2", "PIP3", "Erk", "Akt", "PKA", "PKC", "P38", "Jnk".
+Your task is to determine the causal relationship between the given variables.
+If there are both indirect and direct relationships between the variables, you should describe only the direct one.’
+We repeat the process $10$ times and use the most answered feedback. As its uncertainty, we consider the simple ratio of the number of times the LLM gave that feedback over $10$.
+D.1 Elicitation Framework
+When considering an LLM as an expert, we extend our elicitation framework to accommodate a query-dependent confidence level.
+To understand this, recall that our refinement algorithm iteratively collects a feedback $f_{r}\in\{1,2,3,4\}$ regarding a selected edge $r$ from an expert with confidence $\pi$.
+In particular, our model in Section 4 assumes that $\pi$ does not depend on $r$.
+To relax this constraint, we allow for a $r$-dependent likelihood function in the hierarchical Bayesian model defining our expert-in-the-loop process.
+That is, our model becomes
+$\omega_{r}$ $\sim\text{Cat}(\boldsymbol{\rho}_{r}),$ (29)
+$f_{r}|\omega_{r}$ $\sim\text{Cat}\left(\delta_{\omega_{r}}\cdot\pi_{r}+(\mathbf{1}-\delta_{\omega_{r}})\cdot\left(\frac{1-\pi_{r}}{3}\right)\right)$ (30)
+(using the same notation as in Section 4).
+Although demonstrably effective (see our experiments in Section 5), this approach deviates our model from a purely Bayesian paradigm.
+We leave the investigation of a fully Bayesian approach for knowledge elicitation in causal discovery to future endeavors.
+For example, this could be achieved through the introduction of a Beta prior distribution over $\pi$.
+Also, as explained in the previous section, we elicit $\pi_{r}$ from the LLM through a bootstrapping algorithm.
+In doing so, we draw on recent studies on the calibration of the verbalized confidence of LLMs (Zhang et al.,, 2024; Xiong et al.,, 2024; Yang et al.,, 2024; Tao et al.,, 2024).
+Xiong et al., (2024), for example, shows that LLMs often overstate their confidence and that averaging the LLM’s responses across multiple queries can provide more accurate uncertainty estimates.
+Interestingly, our early (unreported) experiments using the LLM’s verbalized confidence were also less successful than the ones that relied on bootstrapping.
+Appendix E Related Work
+E.1 CD under latent confounding.
+Following the seminal works by Spirtes et al., (2001) and Zhang, 2008b introducing the complete FCI, a variety of works have emerged.
+Among them are algorithms designed for sparse scenarios, including RFCI (Colombo et al.,, 2012) and others (Silva,, 2013; Claassen et al.,, 2013).
+Notably, Silva, (2013)’s framework uses a Bayesian approach to CD of Gaussian causal diagrams based on sparse covariance matrices.
+Nonetheless, it requires sampling one edge at a time and relies on numerical heuristics that might effectively alter the posterior we are sampling from.
+Colombo et al., (2012) introduced the conservative FCI to handle conflicts arising from statistical errors in scenarios with limited data, even though it yields less informative results.
+Subsequent efforts to improve reliability led to the emergence of constraint-based CD algorithms based in Boolean satisfiability (Hyttinen et al.,, 2014; Magliacane et al.,, 2016), although they are known to scale poorly on $|\mathbf{V}|$ (Lu et al.,, 2021).
+In another paradigm, score-based search algorithms rank MAGs according to goodness-of-fit measures, commonly using BIC for linear Gaussian SCMs (Triantafillou and Tsamardinos,, 2016; Zhalama et al., 2017a, ; Rantanen et al.,, 2021).
+There are also hybrid approaches that combine constraint-based strategies to reduce the search space, such as GFCI (Ogarrio et al.,, 2016), M3HC (Tsirlis et al.,, 2018), BCCD (Claassen and Heskes,, 2012) and GSPo (Bernstein et al.,, 2020).
+Continuous optimization has recently emerged as a novel approach to score-based CD, as DCD (Bhattacharya et al.,, 2021) and N-ADMG (Ashman et al.,, 2023).
+While N-ADMG focuses on a more restricted setting compared to AGs, it offers some uncertainty quantification in the variational posterior, making it the most directly comparable baseline.
+We rigorously follow the protocols outlined in the original works.
+E.2 CD with expert knowledge.
+Previous works on CD have explored various forms of background knowledge.
+This includes knowledge on edge existence/non-existence (Meek,, 1995), ancestral constraints (Ribeiro et al.,, 2024; Chen et al.,, 2016), variable grouping (Parviainen and Kaski,, 2017), partial order (Andrews,, 2020) and typing of variables (Brouillard et al.,, 2022).
+Incorporating expert knowledge is pivotal to reducing the search space and the size of the learned equivalence class.
+Nevertheless, due to significant challenges, up to date, there are only a few works trying to integrate background knowledge into CD within the context of latent confounding (Andrews,, 2020; Wang et al.,, 2022).
+These works operate under the assumption of perfect expert feedback.
+In contrast, our contribution is the first to deal with more realistic situations where expert input might be inaccurate.
+Recently, many works have been using the domain knowledge within LLMs to improve CD algorithms.
+Usually, the frameworks focus on using the LLM alongside a data-driven model.
+To achieve this, specialized querying strategies are often employed (Wan et al.,, 2024), for example,
+  1. 1.
+    Pair-wise discovery;
+  2. 2.
+    Conditional independence tests;
+  3. 3.
+    Full graph discovery.
+Most works elicit knowledge from the LLM in pairs of variables (Ban et al.,, 2025; Jiralerspong et al.,, 2024).
+However, there are some that focus on discovering the full graph at once (Ban et al.,, 2023).
+On top of that, recent works have also used LLMs for providing indirect information that improves the CD process.
+For example, i) Li et al., (2025) uses the LLM for intervention targeting, while ii) Vashishtha et al., (2023) solely focus on finding a causal order, instead of the causal graph.
+Remarkably, none of the mentioned works considers any uncertainty metric for the answers given by the LLM, which we do.
+A recent paper by Ankan and Textor, (2025) also considered an expert-in-the-loop which could be either a human or an LLM.
+The approach is based on using CI statements or an iteration of the classical PC algorithm (Spirtes et al.,, 1993) together with an expert-in-the-loop.
+Differently from our work, this outputs a single DAG and does not consider any uncertainty over the final result.
+Moreover, there are experiments using LLMs as experts.
+Importantly, the used dataset does not have a defined true graph to be compared to and the code is not available.
+So, we were unable to test their framework and use it as comparison.
+E.3 Generative Flow Networks
+GFlowNets (Bengio et al., 2021a, ) are generative models that sample discrete composite objects from an unnormalized reward function.
+They have been successfully used to a variety of structures, including protein sequences (Jain et al.,, 2022), molecules (Bengio et al., 2021a, ), schedules (Zhang et al.,, 2023), phylogenetic trees (Anonymous,, 2023), strategies in adversarial games (Jiralerspong et al.,, 2023), dropout masks (Liu and et al.,, 2023), sentences (Hu et al., 2023a, ).
+They have also been used to train energy-based models (Zhang et al.,, 2022; Hu et al., 2023b, ).
+Recently, Garipov et al., (2023) presented a strategy for reusing expensively pretrained GFlowNets based on composition and classifier guidance.
+In the field of structure learning, they have been applied to Bayesian networks, more specifically to sample a posterior over DAGs, although without accounting for unobserved confounding (Deleu et al.,, 2022), and to dynamic Bayesian networks (Atanackovic et al.,, 2023).
+Recently, Deleu et al., (2023) proposed an extension to jointly infer the structure and parameters, also grounded in the assumption of causal sufficiency.
+It is worth highlighting that training GFlowNets in these scenarios presents optimization challenges, resulting in the utilization of a variety of loss functions (Shen et al.,, 2023; Malkin et al.,, 2022; Pan et al., 2023a, ; Pan et al., 2023b, ).
+Moreover, Lahlou et al., (2023) extended GFlowNets to continuous domains.
+Appendix F Details on (Non-)Ancestral Knowledge
+Our EITL approach builds on the premise that experts can determine whether a variable is a cause (ancestor) of another, regardless of whether this causation is direct or confounded.
+This knowledge is much coarser compared to that needed for a causal DAG.
+While directed edges in a causal DAG represent direct causation (assumed to be unconfounded), which can be challenging for experts to identify, edges in an AG represent only the existence of a directed (causal) path, regardless of any other potentially unobserved, confounding paths.
+Scientists often have an understanding of the ancestral (or non-ancestral) relationship between two variables.
+Illustratively, in biomedicine, causal knowledge primarily arises from randomized experiments, readily available on platforms like PubMed (Wheeler et al.,, 2007), DrugBank (Knox et al.,, 2024), and GO-CAM (Thomas et al.,, 2019).
+Yet, it is important to note that causality from experimental studies only guarantees ancestrality, not directness or unconfoundedness.
+For example, causal effects may be mediated by unobserved factors, and confounding, albeit reduced by randomization in experiments, can persist in observational studies.
+An expert can incorporate these established ancestral relations using our framework by increasing the probability of the corresponding directed edges.
+There are also numerous examples of known non-ancestral relationships.
+For instance, it is widely accepted that sociodemographic variables (e.g., age and sex), as well as genetic variables (referred to as $X$), are not caused by factors such as drugs, diseases, or other phenotypes (referred to as $Y$).
+To reflect this understanding that $Y$ is not an ancestor of $X$, one can reduce the likelihood of $Y$ being ancestral to $X$, achieved by decreasing the probability of the corresponding edges $Y\rightarrow X$.
+Additionally, information indicating that an association is purely spurious, with no causation in any direction, may also be accessible.
+For instance, the link between coffee consumption ($X$) and heart disease ($Y$) is likely to be solely spurious, as factors such as smoking, diet, and lifestyle habits among heavy coffee drinkers could confound the results.
+This knowledge is incorporated by increasing the probability of a bidirected edge $X\leftrightarrow Y$.
+Finally, our expert-in-the-loop approach considers the confidence level of such knowledge, facilitating its inclusion even when the source may lack reliability.